### v1

# thanks: Corresponding author

Abstract
In recent years, jumps in the particle circular-orbit radius $r_{\rm c}$ have been proposed as observational signatures of black hole phase transitions, supported by the general consensus that $r_{\rm c}$ increases monotonically with the horizon radius $r_{h}$. However, in this paper, we will show that this monotonicity is not universally valid. By deriving a universal, ensemble-independent criterion for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ and applying it to the $d$-dimensional Reissner-Nordström-Anti-de Sitter black hole, we reveal a sharp dichotomy: while $r_{\rm c}(r_{h})$ exhibits strictly monotonicity along isobars, it typically becomes non-monotonic along isotherms (except for the photon sphere of Schwarzschild-Anti-de Sitter black hole)–a behavior that has been overlooked in previous studies. Nevertheless, we rigorously prove that within the phase-transition region itself, the monotonicity of $r_{\rm c}(r_{h})$ is preserved. However, the generalizability of this result to other spacetimes requires case-by-case verification, and its validity could have notable implications for the observability of the jump in $r_{\rm c}$. Interestingly, non-monotonicity itself acts as a new observational probe: the monotonicity contrast in $r_{\rm c}(r_{h})$ distinguishes isobaric from isothermal evolution, providing a complementary thermodynamic probe. Furthermore, we show that the critical exponents extracted from $r_{\rm c}$ match those obtained from $r_{h}$, independent of ensemble. Remarkably, all these phenomena are closely tied to the thermodynamic first law.
I Introduction
Black hole thermodynamics has evolved into a central theme of gravitational physics. The Hawking-Page transition first revealed a first-order transition between thermal Anti-de Sitter (AdS) radiation and large AdS black holes [1]. Subsequent work has uncovered thermodynamic phase transitions in various black-hole spacetimes, including Van der Waals-type small/large transitions in AdS black holes [2, 3, 4, 5, 6, 7], thereby establishing a rich thermodynamic phase structure.
Yet direct observation of these thermodynamic phenomena remains limited. In contrast, optical and dynamical features–such as the photon sphere and the innermost stable circular orbit (ISCO)–are directly observable and encode essential geometric information relevant to shadows and accretion physics [8, 9, 10]. The EHT image of M87* made this explicit, revealing an emission ring whose boundary traces the photon orbit [11].
This naturally raises the possibility that black hole thermodynamics could be probed through optical features. Compelling evidence comes from the work of Wei et al., who showed that in charged AdS black holes, both the photon-sphere radius and the minimum impact parameter exhibit Van derWaals-type non-monotonicity and jump discontinuities, behaving as mean-field order parameters [12]. Subsequently, related signatures involving photon orbits and ISCOs have been identified across a variety of spacetimes and modified gravitational theories [13, 14, 15].
A central consideration in these studies is the relationship between the circular-orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$. Across a first-order small/large transition, $r_{h}$ undergoes a discontinuous jumps, and the monotonicity $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly–making it a natural geometric order parameter (see Appendix A for a brief mathematical derivation). This expectation was later formalized by Zhang et al., who proved monotonic growth of both radii (the photon-sphere radius and ISCO radius) with $r_{h}$ in general static spherical geometries–a result that has been widely applied in subsequent work [16].
However, we find that once more than one thermodynamic parameter is allowed to vary, this monotonicity claim fails. To illustrate this, we present a simple counterexample based on the Schwarzschild-AdS (Sch-AdS) black hole: the photon-sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and $P$, where $P=-\Lambda/(8\pi)$ with $\Lambda$ being the negative cosmological constant. Increasing $P$ decreases $r_{h}$, with $r_{\rm ps}$ unchanged. Consequently, there exist smooth paths in $(M,P)$ parameter space along which $r_{h}$ increases and $r_{\rm ps}$ decreases. Figure 1 provides an explicit example, and Appendix B presents a fully analytic realization. This breakdown of monotonicity in $r_{\rm c}(r_{h})$ has been neglected in previous studies. The reason is straightforward: most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically grow monotonically with $M$. Even in studies that employed the isothermal ensembles (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not systematically examined, leading to the omission of non-monotonic behavior. Thus, a genuine understanding of how circular orbits encode thermodynamic information requires analyzing the monotonicity of $r_{\rm c}(r_{h})$ along arbitrary parameter paths–precisely the general framework developed in this work.
Figure 1: Behavior of $M(r_{h},P)$ in the Sch-AdS spacetime. Thin black solid curves denote contours of constant mass $M$. Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon-sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
Our strategy is as follows. In Sec. II, we review the conditions for photon spheres and ISCOs in a general static, spherically symmetric spacetime, and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary parameter path. In Sec. III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström-AdS (RN-Ads) black holes, focusing on the isobaric ($T$-$S$) and isothermal ($P$-$V$) thermodynamic paths. In Sec. IV, we show that the critical exponents remain invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, demonstrating that $r_{\rm c}$ defines the same universality class as $r_{h}$. Finally, Sec. V summarizes and discusses our main results.
II Monotonicity criterion
II.1 Review of circular orbits
Consider a static, spherically symmetric spacetime with the line element
$\mathrm{d}s^{2}=-f(r;\lambda)\,\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{f(r;\lambda)}+r^{2}\mathrm{d}\Omega^{2},$ (1)
where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black-hole parameters, such as mass $M$, charge $q$, and so on. The Lagrangian of a particle moving in the spacetime determined by Eq. (1) reads
$\mathcal{L}=\tfrac{1}{2}\,g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=-\tfrac{1}{2}\,\epsilon,\quad\epsilon=\begin{cases}0,&\text{massless particle (photon)},\\ 1,&\text{massive particle}.\end{cases}$ (2)
Where the overdot denotes differentiation with respect to the affine parameter. Without loss of generality, we set $\theta=\pi/2$, restricting the motion to the equatorial plane. Then, the Lagrangian reduces to
$\mathcal{L}=\frac{1}{2}\!\left[-f(r;\lambda)\,\dot{t}^{\,2}+\frac{\dot{r}^{\,2}}{f(r;\lambda)}+r^{2}\dot{\phi}^{\,2}\right]=-\tfrac{1}{2}\,\epsilon.$ (3)
Since the metric is stationary and axisymmetric, the corresponding cyclic coordinates $t$ and $\phi$ give rise to two conserved quantities,
$E:=-\frac{\partial\mathcal{L}}{\partial\dot{t}}=f(r;\lambda)\,\dot{t},\qquad L:=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=r^{2}\dot{\phi},$ (4)
which represent the particle’s energy and angular momentum, respectively. Substituting Eq. (4) into Eq. (3) yields the following radial equation of motion
$\dot{r}^{\,2}=E^{2}-f(r)\!\left(\epsilon+\frac{L^{2}}{r^{2}}\right)\equiv\mathcal{R}(r).$ (5)
The particle’s circular orbits at $r=r_{c}$ are determined by
$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0.$ (6)
Solving Eq. (6) for a massless particle with $\epsilon=0$ gives
$\Phi_{0}(r;\lambda):=r\,\partial_{r}f(r;\lambda)-2f(r;\lambda)=0,\quad\text{at}\,r=r_{\rm ps},$ (7)
where $r_{\rm ps}$ denotes the photon-sphere radius. The condition for an unstable photon sphere further requires
$\mathcal{R}^{\prime\prime}(r_{\rm ps})<0,$ (8)
which is equivalent to
$\Gamma_{0}(\lambda):=\left.\partial_{r}\Phi_{0}(r;\lambda)\right|_{r_{\rm ps}}<0.$ (9)
The ISCO of a massive particle with $\epsilon=1$ marks the transition between stable and unstable circular motion and is defined by the marginal stability condition
$\mathcal{R}(r_{\rm isco})=\mathcal{R}^{\prime}(r_{\rm isco})=\mathcal{R}^{\prime\prime}(r_{\rm isco})=0,$ (10)
which can be recast into a universal form,
$\Phi_{1}(r;\lambda)$ $=r\,f(r;\lambda)\,\partial_{r}^{2}f(r;\lambda)-2r\big(\partial_{r}f(r;\lambda)\big)^{2}$ (11)
$+3f(r;\lambda)\,\partial_{r}f(r;\lambda)=0,\quad\text{at}\,r=r_{\rm isco}.$
To ensure that circular orbits are stable outside ISCO but unstable inside, one further requires its radius $r_{\rm isco}$ to satisfy
$\Gamma_{1}(\lambda):=\left.\partial_{r}\Phi_{1}(r;\lambda)\right|_{r_{\rm isco}}>0.$ (12)
Summarizing the above conditions, we have
$\Phi_{\epsilon}(r;\lambda)=0,\quad\text{at}\,r=r_{c}(\epsilon;\lambda):=\begin{cases}r_{\rm ps}(\lambda),&\epsilon=0,\\ r_{\rm isco}(\lambda),&\epsilon=1,\end{cases}$ (13)
and the corresponding stability criteria
$\Gamma_{\epsilon}(\lambda)\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
II.2 Derivation of the Monotonicity Criterion
To investigate the monotonic relation between the circular-orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
$\Phi_{\epsilon}(r_{c};\lambda)=0,\qquad f(r_{h};\lambda)=0.$ (15)
According to the implicit function theorem, if
$\left.\partial_{r}\Phi_{\epsilon}\right|_{(r_{\mathrm{c}};\lambda)}\neq 0,\qquad\left.\partial_{r}f\right|_{(r_{h};\lambda)}\neq 0,$ (16)
then there exist locally smooth functions $r_{\mathrm{c}}(\lambda)$ and $r_{h}(\lambda)$ in the neighborhoods of $(r_{\mathrm{c}};\lambda)$ and $(r_{h};\lambda)$, respectively. Fortunately, according to Eqs. (9) and (12), one finds
$\left.\partial_{r}\Phi_{\epsilon}\right|_{(r_{\mathrm{c}};\lambda)}=\Gamma_{\epsilon}\neq 0,$ (17)
and
$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa>0,$ (18)
where $\kappa=\tfrac{1}{2}\,\partial_{r}f(r;\lambda)\big|_{(r_{h};\lambda)}$ is the surface gravity of the black hole, which is positive for any nonextremal case. Hence, the conditions (16) are automatically satisfied.
Based on this, Eq. (15) can in principle be differentiated with respect to $\lambda$. Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(t)$ in the parameter space, along which all quantities become functions of a single real variable $t$. This allows us to directly apply the chain rule, yielding
$\frac{\mathrm{d}}{\mathrm{d}t}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(t);\lambda(t)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
$\frac{\mathrm{d}}{\mathrm{d}t}\,f\big(r_{h}(t);\lambda(t)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
where the overdot denotes differentiation with respect to $t$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$. Combining Eqs. (19) and (20), we finally obtain
$\frac{\mathrm{d}r_{\mathrm{c}}}{\mathrm{d}r_{h}}=\frac{\dot{r}_{\mathrm{c}}}{\dot{r}_{h}}=\frac{2\kappa}{\Gamma_{\epsilon}}\,\frac{\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle}{\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle},\qquad(\dot{r}_{h}\neq 0),$ (21)
where we have written $\Phi_{\epsilon}$ and $f$ as shorthand for $\Phi_{\epsilon}(r_{\mathrm{c}};\lambda)$ and $f(r_{h};\lambda)$, respectively, and this convention will be used throughout the paper.
The criterion (21) shows that the monotonic relation between $r_{\mathrm{c}}$ and $r_{h}$ is governed by two independent factors. The first factor, $2\kappa/\Gamma_{\epsilon}$, is purely geometric, determined by the intrinsic properties of the black hole and its circular orbits. As discussed earlier, it remains nonvanishing and therefore does not affect the monotonicity of $r_{\mathrm{c}}(r_{h})$. The second factor, $\langle\nabla_{\lambda}\Phi_{\epsilon},\dot{\lambda}\rangle/\langle\nabla_{\lambda}f,\dot{\lambda}\rangle$, encodes how the parameters evolve along a chosen path in the thermodynamic space. It therefore determines whether the overall monotonicity of $r_{\mathrm{c}}(r_{h})$ is preserved or violated, emphasizing the importance of considering multiparameter constraints in black hole thermodynamics, as discussed in the Introduction.
In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN-AdS black hole.
III Examples of $d$-Dimensional Black Holes under Isobaric and Isothermal Constraints
In $d$-dimensional spacetime, a static and spherically symmetric AdS black hole is described by the line element
$\mathrm{d}s^{2}=-f(r)\,\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{f(r)}+r^{2}\,\mathrm{d}\Omega_{d-2}^{2},$ (22)
where $\mathrm{d}\Omega_{d-2}^{2}$ denotes the line element of a unit (d2)(d-2)-sphere. The metric function takes the general form
$f(r)=1-\frac{c_{d}M}{r^{d-3}}+\frac{q^{2}}{r^{2(d-3)}}+\frac{16\pi P}{(d-1)(d-2)}\,r^{2},$ (23)
where $M$ is the ADM mass, $q$ is the charge parameter, and $P=-\Lambda/(8\pi)$ represents the thermodynamic pressure associated with the cosmological constant $\Lambda$. The coefficients
$c_{d}=\frac{16\pi}{(d-2)\Omega_{d-2}},\qquad\Omega_{d-2}=\frac{2\pi^{(d-1)/2}}{\Gamma((d-1)/2)},$ (24)
depend only on the spacetime dimension, with $\Omega_{d-2}$ being the area of the unit (d2)(d-2)-sphere. The metric (23) reproduces several well-known limits:
  • $q=0$: the $d$-dimensional Sch-AdS black hole;
  • $q\neq 0$: the RN-AdS black hole;
  • $P\to 0$: the asymptotically flat Schwarzschild or RN solution in $d$ dimensions.
III.1 Isobaric Process
The isobaric process corresponds to the $T$-$S$ ensemble commonly employed in black hole thermodynamics. Here, all thermodynamic variables except the mass remain fixed. Thus, we may take $\lambda=(M;\lambda_{\mathrm{fixed}})$ and choose $M$ itself as the path parameter. This leads to $\dot{r}_{h}={\mathrm{d}r_{h}}/{\mathrm{d}M}$. To examine whether $\dot{r}_{h}$ can vanish, we consider the thermodynamic first law
$\mathrm{d}M=T\,\mathrm{d}S\quad\Rightarrow\quad\frac{\mathrm{d}r_{h}}{\mathrm{d}M}=\frac{1}{T\,S^{\prime}(r_{h})},$ (25)
where $T={2\kappa}/{4\pi}=\left.{\partial_{r}f(r;\lambda)}/{4\pi}\right|_{(r_{h};\lambda)}>0$ is the Hawking temperature, and $S^{\prime}(r_{h})=\mathrm{d}S/\mathrm{d}r_{h}$. In $d$-dimensional general relativity, S=A/4S={A}/{4} with $A=\Omega_{d-2}r_{h}^{\,d-2}>0$, so that
$S^{\prime}(r_{h})=\frac{\Omega_{d-2}}{4}\,(d-2)\,r_{h}^{\,d-3}>0\quad\Rightarrow\quad\frac{\mathrm{d}r_{h}}{\mathrm{d}M}>0.$ (26)
Hence, taking $M$ as the path parameter is fully justified, and the general criterion (21) simplifies to
$\frac{\mathrm{d}r_{\mathrm{c}}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}.$ (27)
The monotonicity of $r_{\rm c}(r_{h})$ is therefore governed by $\partial_{M}\Phi_{\epsilon}$. Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ determines the background geometry and hence controls the response of circular orbits, which generically vary with $M$.
For the metric (23), we obtain
$\partial_{M}\Phi_{0}$ $=(d-1)\frac{c_{d}}{r_{\rm ps}^{\,d-3}}>0,$ (28)
$\partial_{M}\Phi_{1}$ $=-\frac{2\,r_{\rm isco}^{\,1-4d}}{(d-1)(d-2)}\,\sqrt{h(r_{\rm isco},q,P,d)}<0,$
where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into Eq. (27) and using Eqs. (14) and (26), we find
$\epsilon=0:$ $\quad-\frac{1}{\Gamma_{0}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}>0,\ \partial_{M}\Phi_{0}>0\ \Rightarrow\ \frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}>0,$ (29)
$\epsilon=1:$ $\quad-\frac{1}{\Gamma_{1}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<0,\ \partial_{M}\Phi_{1}<0\ \Rightarrow\ \frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}>0.$ (30)
Hence, along an isobaric process, both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ increase monotonically. The opposite signs of $\Gamma_{0}$ and $\Gamma_{1}$ are exactly compensated by those of $\partial_{M}\Phi_{0}$ and $\partial_{M}\Phi_{1}$, resulting in a consistent monotonic relation between $r_{\rm c}$ and $r_{h}$. This implies that, during an isobaric process, a small-to-large black hole phase transition necessarily induces a corresponding discontinuous jump in both $r_{\mathrm{ps}}$ and $r_{\mathrm{isco}}$. Such abrupt changes could, in principle, be detected through astronomical observations, providing a potential signature of black hole phase transitions. These findings are consistent with the earlier results reported by Wei, Zhang, et al.
Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional (4D) RN-AdS black hole, one obtains
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{2\kappa}{{c_{d}}/{r_{h}^{\,d-3}}}=\frac{2(-{q^{2}}/{r_{h}^{3}}+{1}/{r_{h}}+8\pi P\,r_{h})}{{c_{d}}/{r_{h}^{\,d-3}}}>0,$
$\Gamma_{0}$ $=\frac{4q^{2}-2r_{\rm ps}^{2}}{r_{\rm ps}^{3}}<0,$
$\Gamma_{1}$ $=\frac{1}{6\,r_{\rm isco}^{6}}\Bigl[15q^{4}+r_{\rm isco}^{4}+352\pi P\,r_{\rm isco}^{6}+4800\pi^{2}P^{2}r_{\rm isco}^{8}$
$\quad+\,r_{\rm isco}^{2}\sqrt{g}-120\pi P\,r_{\rm isco}^{4}\sqrt{g}$
$\quad+\,q^{2}\!\left(48\pi P\,r_{\rm isco}^{4}-9\sqrt{g}\right)\Bigr]>0,$ (31)
where
[…]
It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq. (25). Using Eq. (27) together with the above results, we find
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3\,r_{\rm ps}^{2}\!\left(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4}\right)}{2\left(2q^{2}-r_{\rm ps}^{2}\right)r_{h}^{2}},$ (33)
and
$\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=-\,6\,(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4})\,r_{\rm isco}^{13}\,\sqrt{g}\;/\\ \Bigl[r_{h}^{2}\bigl(15q^{4}r_{\rm isco}^{11}+r_{\rm isco}^{15}+352\pi P\,r_{\rm isco}^{17}+4800\pi^{2}P^{2}r_{\rm isco}^{19}\\ +48\pi P\,q^{2}\,r_{\rm isco}^{15}+(r_{\rm isco}^{2}-120\pi P\,r_{\rm isco}^{4}-9q^{2})\,r_{\rm isco}^{11}\,\sqrt{g}\bigr)\Bigr].$ (34)
In the Sch-AdS limit ($q=0$), the above relations reduce to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2}+12\pi Pr_{h}^{2},$ (35)
[…]
where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further taking the Schwarzschild limit ($q=0$, $P=0$), one obtains
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2},\qquad\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=3,$ (37)
which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$.
In Figs. 2 and 3, we present the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$ and $r_{\mathrm{isco}}(r_{h})$ for the four-dimensional case. As shown, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch-AdS black hole, whereas for the RN-AdS black hole it first decreases and then increases, with the extremum occurring precisely at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$. Truncating the curve at this point, $r_{\mathrm{c}}(r_{h})$ retains its local monotonicity. These numerical results are in complete agreement with Eqs. (33)–(36).
Figure 2: Behavior of $r_{\rm ps}(r_{h})$ along isobaric paths in $d=4$. For the Sch-AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.
Figure 3: Behavior of $r_{\rm isco}(r_{h})$ along isobaric paths in $d=4$
III.2 Isothermal Process
The isothermal process corresponds to the $P$-$V$ ensemble commonly employed in black hole thermodynamics. Here the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$. Therefore, we take $\lambda=(M,P;\lambda_{\mathrm{fixed}})$. Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter gives (see Appendix C for justification)
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right).$ (38)
In general, this relation has no simple closed form. However, for photons ($\epsilon=0$), many black-hole metrics satisfy $\partial_{P}\Phi_{0}=0$, for example when the pressure enters the metric only through a term proportional to $Pr^{2}$. In such cases, the criterion (38) reduces to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{0}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\partial_{M}\Phi_{0}.$ (39)
This result is consistent with the criterion for the isobaric process, but differs in that under the isobaric constraint the first law of thermodynamics requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}\neq 0$, ensuring that $r_{\mathrm{ps}}(r_{h})$ remains monotonic. In contrast, under the isothermal constraint the first law takes the form
$\mathrm{d}M=T_{0}\,\mathrm{d}S+V\,\mathrm{d}P,$ (40)
where $V$ is the thermodynamic volume. Dividing both sides by $\mathrm{d}r_{h}$ yields
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (41)
indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become non-monotonic. This behavior is physically intuitive: when ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$, one must have
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\frac{T_{0}}{V}\,S^{\prime}(r_{h})<0.$ (42)
This means that, near an extremum of $M(r_{h})$, the same $M$ corresponds to two distinct $P$. As $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, consequently a single $r_{\rm ps}$ maps to two different $r_{h}$, and $r_{\rm ps}(r_{h})$ is necessarily non-monotonic. Next, we apply the criterion (38) to the metric (23).
III.2.1 Photon sphere ($\epsilon=0$)
For the metric (23) one obtains
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{1}{(d-1)\,c_{d}}\Big[$  $2(d-3)\,r_{h}^{\,d-4}+4(d-2)\,\pi T_{0}\,r_{h}^{\,d-3}$ (43)
$-2(d-2)(d-3)\,q^{2}\,r_{h}^{\,2-d}\Big],$
and
$\Gamma_{0}=\frac{2(d-3)}{r_{\rm ps}}\Big[(d-2)\,q^{2}\,r_{\rm ps}^{\,6-2d}-1\Big].$ (44)
Moreover, $\partial_{P}\Phi_{0}=0$, and $\partial_{M}\Phi_{0}$ agrees with Eq. (28).
We now analyze the sign of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$. For $q=0$ one immediately has ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$. For $q\neq 0$, differentiating ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ gives
$\frac{\mathrm{d}^{2}M}{\mathrm{d}r_{h}^{2}}=\frac{1}{(d-1)c_{d}}$ [ $\Big[2(d-3)(d-4)\,r_{h}^{\,d-5}$ (45)
$+4(d-2)(d-3)\pi T_{0}\,r_{h}^{\,d-4}$
$+2(d-2)^{2}(d-3)q^{2}\,r_{h}^{\,1-d}\Big]>0,$
so ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is strictly increasing in $r_{h}$. Meanwhile, From Eq. (43),
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (46)
hence there exists a unique $r_{h0}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h0}}=0$. By the criterion (39), it follows that:
  • $q=0$: $r_{\rm ps}(r_{h})$ is monotonically increasing;
  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is non-monotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h0}$, and then increases.
This result reveals that the presence of electric charge $q$ breaks the monotonic behavior of $r_{\mathrm{ps}}(r_{h})$, making it globally non-monotonic–an aspect that has been omitted in previous studies. The explicit expression for ${dr_{\mathrm{ps}}}/{dr_{h}}$ reads
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}$ $=\Big[(d-3)(2-d)\,q^{2}\,r_{h}^{6}$ (47)
$\qquad+\big(d-3+2(d-2)\pi T_{0}\,r_{h}\big)r_{h}^{2d}\Big]\,r_{\rm ps}^{\,4+d}\;/$
$\quad\Big[(d-3)\big((2-d)\,q^{2}\,r_{\rm ps}^{6}+r_{\rm ps}^{2d}\big)\,r_{h}^{\,4+d}\Big],$
For $d=4$ (the RN-AdS case), this reduces to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{r_{\rm ps}^{2}\,\left[2q^{2}-r_{h}^{2}\,(1+4\pi T_{0}\,r_{h})\right]}{(2q^{2}-r_{\rm ps}^{2})\,r_{h}^{2}},$ (48)
and, in the Sch-AdS limit ($q=0$), further simplifies to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=1+4\pi T_{0}\,r_{h}.$ (49)
This coincides with the standard Sch-AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
Figure 4 presents the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$. For the Sch-AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$ and increases monotonically thereafter, remaining globally monotonic in full agreement with Eq. (49). In contrast, the RN-AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
The breakdown of monotonicity naturally raises an important question: does it affect the discontinuous jump of $r_{\mathrm{ps}}$ across the phase transition, potentially obscuring its optical signature?
The answer is: possibly. If the small and large black hole radii, $r_{h,s}$ and $r_{h,l}$, lie on opposite sides of the extremum of $r_{\mathrm{ps}}(r_{h})$, the amplitude of the discontinuous change in $r_{\mathrm{ps}}$ may be reduced, requiring higher observational resolution to detect it. In the extreme case where $r_{\mathrm{ps}}(r_{h,s})=r_{\mathrm{ps}}(r_{h,l})$, the discontinuity would vanish entirely, rendering the phase transition optically unobservable.
From a thermodynamic perspective, however, this is not necessarily a drawback. As shown earlier, $r_{\mathrm{c}}(r_{h})$ remains monotonic in the isobaric process, while $r_{\mathrm{ps}}(r_{h})$ becomes non-monotonic in the isothermal one. Such a contrast in optical behavior between the two ensembles offers a potential diagnostic for identifying the underlying thermodynamic process of a black hole– a perspective that provides a new avenue for probing black hole thermodynamics, to be explored in future work.
Another possibility is that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{ps}}(r_{h})$, so that $r_{\mathrm{ps}}(r_{h})$ remains locally monotonic throughout the phase-transition region. This scenario indeed occurs for the RN-AdS black hole studied here, and in Appendix F we provide an explicit proof for the $d=4$ case.
Figure 5 shows the numerically computed profiles of $P(r_{h})$ and $P(r_{\mathrm{ps}})$ at fixed parameters, where the characteristic S-shaped behavior of the phase transition is fully preserved. In this configuration, one can observe both the global non-monotonicity of $r_{\mathrm{ps}}(r_{h})$ and the discontinuous jump of $r_{\mathrm{ps}}$ across the transition. The coexistence of these two optical signatures provides an especially powerful means for extracting thermodynamic information from black holes.
Figure 4: Behavior of $r_{\rm ps}(r_{h})$ along isothermal paths in $d=4$. The marked point on the curve corresponds to $P=0$ at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated. For the Sch-AdS case, the temperature is likewise set to the value $T=0.8\,T_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.
Figure 5: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$.
III.2.2 ISCO ($\epsilon=1$)
The metric (23) yields
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\frac{(d-2)}{8\pi}\Bigl[$ $(d-2)(d-3)\,q^{2}\,r_{h}^{\,3-2d}$ (50)
$+\bigl(3-d+2\pi T_{0}r_{h}\bigr)\,r_{h}^{-3}\Bigr],$
and
$\partial_{P}\Phi_{1}=\frac{16\pi\,r_{\mathrm{isco}}}{(d-2)(d-1)}$ [ $\Bigl[8+4(d-1)(d-2)q^{2}r_{\mathrm{isco}}^{6-2d}$ (51)
$-c_{d}(d-1)(d+1)M\,r_{\mathrm{isco}}^{3-d}\Bigr],$
where $\partial_{P}\Phi_{1}>0$ (see Appendix D for proof). Since van der Waals-type behavior occurs only for charged black holes ($q\neq 0$), we analyze the RN-AdS and Sch-AdS ($q=0$) cases separately. Moreover, because the explicit expression for ${\mathrm{d}r_{\mathrm{c}}}/{\mathrm{d}r_{h}}$ is algebraically cumbersome, we focus on whether it changes sign rather than displaying the full formula.
(i) Sch-AdS ($q=0$)
For $q=0$, Eq. (50) reduces to
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\,\frac{(d-2)\,\bigl(3-d+2\pi T_{0}r_{h}\bigr)}{8\pi\,r_{h}^{3}}.$ (52)
It has a single zero at $r_{h1}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h1}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h1}$. Using the criterion (38), together with ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}<0$, $\Gamma_{1}>0$, and $\partial_{P}\Phi_{1}>0$, one thus expects $r_{\rm isco}(r_{h})$ to first decrease and then increase. To make this more explicit, we rewrite the criterion (38) as
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\delta(r_{\rm isco},r_{h},d,q,T_{0}).$ (53)
For sufficiently large $r_{h}$ one has ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, which forces $\delta<0$. It remains to verify that $\delta>0$ for small $r_{h}$, ensuring a single zero crossing. In this spirit we find
$\begin{cases}\delta\bigl(3r_{h_{\min}},\,r_{h_{\min}},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h_{\min}},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (54)
where $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h_{\min}})=0$. For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h_{\min}}$. For $d>4$ and $P\to 0$, the higher-dimensional Schwarzschild black hole has no ISCO, hence $r_{\rm isco}\to+\infty$. Equation (54) then confirms that $\delta$ indeed crosses zero. Consequently, under the isothermal constraint, the function $r_{\rm isco}(r_{h})$ is non-monotonic.
(ii) RN-AdS ($q\neq 0$)
Since the RN-AdS black hole exhibits a van der Waals–like phase transition, one has
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\leq 0\quad\text{for}\,T\geq T_{c},$ (55)
where the equality holds precisely at the critical point $(T_{c},P_{c},r_{hc})$. According to the criterion (38), when $r_{h}\geq r_{h0}$ (where $r_{h0}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically. For $r_{h}<r_{h0}$, however, $\delta$ may become positive. To determine this behavior, we examine the limit $r_{h}\to 0^{+}$.
From Eq. (41), one obtains
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (56)
which substituted into the criterion (38) gives
$\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\!\left(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\right).$ (57)
Meanwhile, from Appendix E, as $r_{h}\to 0^{+}$ one has
$V\,\partial_{M}\Phi_{1}\ \longrightarrow\ -\infty,\qquad\partial_{P}\Phi_{1}\ \longrightarrow\ 0.$ (58)
Hence,
$\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\Bigl(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\Bigr)>0,\quad\Rightarrow\quad\frac{\mathrm{d}r_{\mathrm{isco}}}{\mathrm{d}r_{h}}<0.$ (59)
This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic over the full domain of $r_{h}$.
The same conclusion holds for $T<T_{c}$. Along both the small–black–hole and large–black–hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above. The only difference is that for $T<T_{c}$ an intermediate spinodal (unstable) branch appears, on which $\mathrm{d}P/\mathrm{d}r_{h}>0$. Across this segment, the sign of $\delta$ may reverse, thereby altering the local monotonicity of $r_{\mathrm{isco}}(r_{h})$. In all cases, however, under the isothermal constraint $r_{\mathrm{isco}}(r_{h})$ is globally non-monotonic.
Figure 6 presents the numerically computed profiles of $r_{\mathrm{isco}}(r_{h})$. It can be seen that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for both charged and uncharged cases. The distinction lies in that, for the uncharged (Sch-AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ where $P(r_{h})=0$, whereas for the charged (RN-AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
Appendix F likewise shows that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{isco}}(r_{h})$. Consequently, $P(r_{\mathrm{isco}})$ preserves the full local SS–shaped structure of $P(r_{h})$ across the phase transition. An explicit example for fixed parameters is shown in Fig. 7.
Figure 6: Behavior of $r_{\rm isco}(r_{h})$ along isothermal paths in $d=4$.
Figure 7: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm isco})$ for $d=4$ and $q=1$.
It is worth noting that, according to the first law of thermodynamics, a non-monotonic $M(r_{h})$ implies that the same mass $M$ can correspond to two distinct pressures $P$. Since both $r_{\mathrm{isco}}$ and $r_{h}$ depend on $P$, one might expect $r_{\mathrm{isco}}(r_{h})$ to be less prone to non-monotonic behavior. However, our results show that non-monotonicity indeed occurs, even for the Sch-AdS black hole.
To better understand this, we examine the criterion (38). Because $\partial_{M}\Phi_{1}<0$ and $\partial_{P}\Phi_{1}>0$, $r_{\mathrm{isco}}(r_{h})$ is guaranteed to be monotonic as long as ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ have opposite signs. Yet the first law,
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (60)
requires that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ share the same sign in certain ranges of $r_{h}$.
For the $d$-dimensional RN-AdS black hole, ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is non-monotonic and can take negative values. Since the first law demands ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, the two derivatives become of the same sign, allowing the criterion (38) to cross zero. For the $d$-dimensional Sch-AdS black hole, comparing
$T_{0}S^{\prime}(r_{h})=\frac{4\pi}{c_{d}}\,(d-2)\,T_{0}\,r_{h}^{\,d-3},$ (61)
with
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{2(d-3)\,r_{h}^{\,d-4}}{(d-1)c_{d}}+\frac{4\pi(d-2)T_{0}\,r_{h}^{\,d-3}}{(d-1)c_{d}},$ (62)
immediately reveals their relative size in the small-horizon regime. For instance, at the minimal value $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, one finds
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{d}{d-1}\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3},$ (63)
$T_{0}S^{\prime}(r_{h})$ $=(d-2)\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3}.$ (64)
It is clear that
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<T_{0}S^{\prime}(r_{h}),$
which, through the first law $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ requires $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}>0$. Thus ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ acquire the same sign, allowing the criterion (38) to cross zero.
IV Critical-exponent invariance
Critical exponents quantify the universal scaling near the black-hole critical point. They are conventionally defined with the horizon radius $r_{h}$, yet the interchangeability of $r_{\rm c}$ with $r_{h}$ suggests that one may equally define them using the circular-orbit radius $r_{\rm c}$. In this section we prove that, for both isobaric and isothermal constraints, the mapping $r_{\rm c}(r_{h})$ is locally monotonic at criticality, so the replacement $r_{h}\mapsto r_{\rm c}$ amounts to a smooth linear reparametrization in a neighborhood of the critical point. Consequently, critical exponent defined with $r_{\rm c}$ coincides with its counterpart defined with $r_{h}$.
The order-parameter exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
$\Delta r_{h}:=r_{h,l}-r_{h,s}\sim|t|^{\beta},\qquad t:=\frac{T-T_{c}}{T_{c}}\to 0^{-},$ (65)
with the mean-field value $\beta=1/2$. We will show that the circular-orbit radius inherits the same scaling, namely
$\Delta r_{c}\;:=\;r_{c}\!\big(r_{h,l}\big)-r_{c}\!\big(r_{h,s}\big)\;\sim\;|t|^{1/2}.$ (66)
We first perform a Taylor expansion of $P(r_{h},T)$ about the critical point $(r_{h,c},T_{c})$ up to the first nonvanishing orders allowed by the criticality conditions $\bigl.(\partial_{r_{h}}P)\bigr|_{c}=0$ and $\bigl.(\partial_{r_{h}}^{2}P)\bigr|_{c}=0$:
$P(r_{h},T)-P_{c}$ $=\left.(\partial_{T}P)\right|_{c}\,\Delta T+\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T\,\Delta r$ (67)
$+\frac{1}{2}\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)^{2}$
$+\frac{1}{6}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{3}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
where $\Delta r:=r_{h}-r_{h,c}$ and $\Delta T:=T-T_{c}$. At fixed temperature $T$, differentiating with respect to $r_{h}$ gives
$\partial_{r_{h}}P(r_{h},T)$ $=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T$ (68)
$+\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)$
$+\frac{1}{2}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{2}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
On the coexistence line for $T<T_{c}$, the Eq. (65) gives
$\Delta r=(r_{h,l}-r_{h,c})-(r_{h,s}-r_{h,c})=\Delta r_{h}\sim|t|^{1/2},$ (69)
so each power counts as $\Delta r^{n}\sim|t|^{n/2}$. Substituting this into (67) and keeping the leading orders in $|t|$ yields
$\partial_{r_{h}}P(r_{h},T)$ (70)
$=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,T_{c}\,t+\frac{1}{2}\left.(\partial_{r_{h}}^{3}P)\right|_{c}\,|t|+\mathcal{O}\!\big(|t|^{3/2}\big),$
Thus $\bigl(\partial P/\partial r_{h}\bigr)_{T}=\mathcal{O}(|t|)\to 0$ as $t\to 0^{-}$. Along an isotherm, our criterion
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right)$ (71)
combines with the first law
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ (72)
to give, as $t\to 0^{-}$,
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{c}\,S^{\prime}(r_{h,c})+\mathcal{O}(|t|),\qquad\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=\mathcal{O}(|t|).$ (73)
Since $S^{\prime}(r_{h,c})>0$, $V>0$, and $\Gamma_{\epsilon}\neq 0$, $\partial_{M}\Phi_{\epsilon}$ and $\partial_{P}\Phi_{\epsilon}$ are finite at the critical point, we obtain the fixed-sign leading term
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{T_{c}\,S^{\prime}(r_{h,c})}{\Gamma_{\epsilon}}\,\partial_{M}\Phi_{\epsilon}\ +\ \mathcal{O}(|t|),$ (74)
so $r_{c}(r_{h})$ is strictly monotonic for sufficiently small $|t|$. (In the photon case, $\partial_{P}\Phi_{0}=0$ further collapses the subleading $\mathcal{O}(|t|)$ contribution.)
Monotonicity (74) implies the first-order linearization at $r_{h,c}$,
$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\bigl(|r_{h}-r_{h,c}|\bigr),$ (75)
where $C:=\left.{dr_{c}}/{dr_{h}}\right|_{c}\neq 0$. Therefore on the coexistence line
$\Delta r_{c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (76)
and the critical exponent is unchanged:
$\beta_{(r_{c})}=\beta_{(r_{h})}=\tfrac{1}{2}\quad\text{(mean-field)}.$ (77)
For the isobaric process, $r_{c}(r_{h})$ is globally monotonic, so the above conclusion applies verbatim. In summary, under either isobaric or isothermal constraints, the reparameterization $r_{h}\!\mapsto\!r_{c}$ does not change the critical exponents. This is consistent with the numerical findings of Wei et al. in the isobaric photon-sphere case [12], for which we have provided an analytic proof.
V Discussion
In this work, we derived the general criterion (21), applicable to any parameter path in spherically symmetric black holes, which clarifies how circular orbits encode thermodynamic behavior. The criterion reveals that the monotonic relation between the circular-orbit radius $r_{\mathrm{c}}$ and the horizon radius $r_{h}$ is influenced by two factors: the intrinsic properties of the black hole and its circular orbits, and the parameter path chosen in the extended thermodynamic space. While the former never breaks monotonicity, the latter plays a decisive role. Motivated by this, we investigated two thermodynamically significant paths–the isobaric and isothermal processes–using the $d$-dimensional RN-AdS black hole as a test case.
Our analysis demonstrates that the first law of thermodynamics directly constrains the monotonicity of $r_{\mathrm{c}}(r_{h})$. Under the isobaric constraint, the first law enforces $\mathrm{d}M/\mathrm{d}r_{h}\neq 0$, ensuring that $M(r_{h})$ and hence $r_{\mathrm{c}}(r_{h})$ remain monotonic. In contrast, under the isothermal constraint the first law allows $\mathrm{d}M/\mathrm{d}r_{h}=0$, which decouples the extrema of $M(r_{h})$ and $P(r_{h})$ so that a single mass $M$ corresponds to two distinct pressures $P$. This inevitably makes $r_{\mathrm{ps}}(r_{h})$ non-monotonic, since the photon-sphere radius $r_{\mathrm{ps}}$ is $P$-independent whereas $r_{h}$ depends on $P$. An exception occurs in the Sch-AdS limit ($q=0$), where $M(r_{h})$ remains monotonic; the emergence of electric charge $q$ thus breaks this monotonicity and, in turn, that of $r_{\mathrm{ps}}(r_{h})$.
Interestingly, for the ISCO case, both $r_{\mathrm{isco}}$ and $r_{h}$ depend explicitly on $P$, which might seem to favor monotonicity. However, for the RN-AdS family we find that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for all $d$, including the Sch-AdS limit. This arises because the first law requires $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ to share the same sign in certain domains, making the prefactor in the criterion (38) change sign and thus breaking monotonicity. Overall, our results indicate that $r_{\mathrm{c}}(r_{h})$ tends to be monotonic under isobaric evolution but non-monotonic under isothermal evolution.
It is worth noting that this non-monotonicity does not spoil the monotonic behavior of $r_{\rm c}(r_{h})$ in the local phase-transition region, nor does it affect its behavior near the critical point. As a result, the jump in $r_{\rm c}$ remains a reliable signature of the transition, and the critical exponents defined through $r_{\rm c}$ coincide with those obtained from $r_{h}$. These features are closely tied to the black hole first law and are expected to extend to more general classes of black holes.
These results fill an important gap in the study of the non-monotonic behavior of $r_{\mathrm{c}}(r_{h})$ and provide a more robust foundation for identifying black hole phase transitions through optical observables. They also offer an additional avenue for probing black hole thermodynamics: the monotonicity (or lack thereof) of $r_{\mathrm{c}}(r_{h})$ can be used to distinguish whether a black hole evolves along an isobaric or an isothermal process.
This idea is physically well motivated. By the entropy-increase law, the horizon radius $r_{h}$ must grow with time, and if $r_{\mathrm{c}}(r_{h})$ is non-monotonic, the resulting optical appearance of the black hole will exhibit correspondingly distinct behavior. Moreover, the influence of black hole thermodynamics on the monotonicity of $r_{\mathrm{c}}(r_{h})$ is itself a noteworthy feature deserving further investigation.
Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
Appendix A Preservation of Functional Behavior under Monotonic Reparameterization
Suppose that $F(x)$ is a differentiable function defined on an interval $I\subset\mathbb{R}$, and let $x_{m}$ denote its $m$-th zero, so that $F(x_{m})=0$. Now introduce a new variable $y$ and assume a smooth, strictly monotonic mapping
$x=x(y),\qquad y\in J\subset\mathbb{R},$ (78)
with ${\mathrm{d}x}/{\mathrm{d}y}\neq 0$ for all $y\in J$. We then reparameterize $F(x)$ as
$\widetilde{F}(y)=F\bigl(x(y)\bigr).$ (79)
Applying the chain rule gives
$\frac{\mathrm{d}\widetilde{F}}{\mathrm{d}y}=\frac{\mathrm{d}F}{\mathrm{d}x}\,\frac{\mathrm{d}x}{\mathrm{d}y}.$ (80)
Since $\mathrm{d}x/\mathrm{d}y\neq 0$, one immediately has
$\frac{\mathrm{d}\widetilde{F}}{\mathrm{d}y}=0\quad\Longleftrightarrow\quad\frac{\mathrm{d}F}{\mathrm{d}x}=0.$ (81)
Thus, the stationary points of $F$ and $\widetilde{F}$ correspond one-to-one: each stationary point $x_{m}$ of $F$ maps to a stationary point $y_{m}=x^{-1}(x_{m})$ of $\widetilde{F}$, and vice versa.
To demonstrate that the nature of the stationary points is preserved, we differentiate once more:
$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=F^{\prime\prime}\bigl(x(y)\bigr)\,[x^{\prime}(y)]^{2}+F^{\prime}\bigl(x(y)\bigr)\,x^{\prime\prime}(y).$ (82)
At a stationary point $y_{m}$ where $F^{\prime}\bigl(x(y_{m})\bigr)=0$, this reduces to
$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=F^{\prime\prime}(x_{m})\,[x^{\prime}(y_{m})]^{2}.$ (83)
Since $x^{\prime}(y_{m})]^{2}>0$, it follows that
$F^{\prime\prime}(x_{m})>0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})>0,$ (84)
$F^{\prime\prime}(x_{m})<0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})<0.$
Hence, under any smooth and monotonic reparameterization $x\mapsto y$, a local minimum (maximum) of $F$ remains a local minimum (maximum) of $\widetilde{F}$.
Away from the stationary points, Eq. (80) implies that when $\mathrm{d}x/\mathrm{d}y>0$, the signs of $\mathrm{d}\widetilde{F}/\mathrm{d}y$ and $\mathrm{d}F/\mathrm{d}x$ coincide, whereas for $\mathrm{d}x/\mathrm{d}y<0$ they are opposite. Geometrically, the former corresponds to a horizontal stretching of the original curve, while the latter involves both stretching and a reflection about the $x$-axis. In either case, the overall qualitative profile of the function remains unchanged.
If $F$ is identified with the thermodynamic temperature $T$ or pressure $P$ of a black hole, and the variables are smoothly and monotonically related as $x=r_{h}$ and $y=r_{c}$, then under the mapping $r_{h}\mapsto r_{c}$ the functions $T(r_{c})$ and $P(r_{c})$ retain their characteristic van der Waals–like $S$-shaped structure. As a result, during the small-to-large black-hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
Appendix B Constructive counterexample in Sch-AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
For the four-dimensional Sch-AdS black hole, the metric function reads
$f(r)=1-\frac{2M}{r}+\frac{8\pi P}{3}\,r^{2},$ (85)
where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon-sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
$r_{\rm ps}=3M,$ (86)
indicating that $r_{\rm ps}$ is independent of $P$. The event horizon radius $r_{h}$ satisfies $f(r_{h};M,P)=0$, from which we obtain
$M=\frac{r_{h}}{2}+\frac{4\pi P}{3}\,r_{h}^{3}.$ (87)
Eliminating $M$ between Eqs. (86) and (87) yields
$r_{\rm ps}(r_{h})=\frac{3r_{h}}{2}+4\pi P\,r_{h}^{3}.$ (88)
Figure 8 illustrates $r_{\rm ps}(r_{h})$ for nine representative values of $P$, without imposing any specific relation between $P$ and $M$.
Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch-AdS black hole. Colored solid curves denote nine different pressures $P$, and the black dashed curve shows the path-constrained case with $k<0$ and $b>0$.
Next, we impose the path constraint
$P(M)=-\frac{3k^{2}\,\bigl(b+(-1+2k)M\bigr)}{8\pi\,(b-M)^{3}},$ (89)
where $k$ is a dimensionless constant and $b$ carries dimensions of mass. Substituting Eqs. (89), (87), and (86) gives a simple linear relation,
$r_{\rm ps}(r_{h})=3\,(k\,r_{h}+b).$ (90)
In Fig. 8, this corresponds to the black dashed line, which decreases monotonically between points $A$ and $B$. Its intersections with the colored curves are uniquely determined by the constraint (89). Along this dashed segment, all physical requirements are satisfied: $M>0$, $r_{h}>0$, $P>0$, and the surface gravity remains strictly positive,
$\kappa=\frac{1}{2}\left.\partial_{r}f(r)\right|_{r_{h}}=\frac{M}{r_{h}^{2}}+\frac{8\pi Pr_{h}}{3}>0,$ (91)
ensuring non-extremality.
This simple example illustrates that the monotonicity of $r_{\rm ps}(r_{h})$ [and likewise of $r_{\rm isco}(r_{h})$] is governed not only by the functional dependence of $(M,P)$ in $f(r)$, but also by the parameter path that links $M$ and $P$. Appropriate constraints can thus lead to a non-increasing behavior of $r_{\rm ps}(r_{h})$. Applying the criterion (21) yields
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{2\kappa}{\Gamma_{0}}\,\frac{\partial_{M}\Phi_{0}}{\partial_{M}f+\partial_{P}f\,\frac{\mathrm{d}P}{\mathrm{d}M}},$ (92)
where
$\frac{\mathrm{d}P}{\mathrm{d}M}=-\frac{3k^{2}\,\bigl(b+bk+2kM-M\bigr)}{4\pi\,(b-M)^{4}}.$ (93)
After straightforward algebra, we obtain
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=3k,$ (94)
which agrees with Eq. (90).
Appendix C Local Existence of $M(r_{h})$ and $P(r_{h})$
Consider the two constraint equations at the horizon,
$f(r_{h};M,P)=0,\qquad T(r_{h};M,P)=T_{0},$ (95)
where both $f$ and $T$ are assumed to be smooth in $(M,P)$. Differentiating these relations and treating $r_{h}$ as the path parameter gives
$\partial_{r_{h}}f\,dr_{h}+\partial_{M}f\,dM+\partial_{P}f\,dP$ $=0,$ (96)
$\partial_{r_{h}}T\,dr_{h}+\partial_{M}T\,dM+\partial_{P}T\,dP$ $=0.$
Eliminating $\mathrm{d}P$ and $\mathrm{d}M$ in turn yields
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=-\,\frac{\partial_{r_{h}}f\,\partial_{P}T-\partial_{P}f\,\partial_{r_{h}}T}{\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T},$ (97)
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ $=-\,\frac{\partial_{r_{h}}f\,\partial_{M}T-\partial_{M}f\,\partial_{r_{h}}T}{\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T}.$
Defining
$\Delta:=\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T,$ (98)
we see that the nondegeneracy condition $\Delta\neq 0$ guarantees–by the implicit function theorem–the local existence and uniqueness of smooth functions $M(r_{h})$ and $P(r_{h})$ satisfying the two constraints in a neighborhood of a reference point. When $\Delta=0$, this parametrization in terms of $r_{h}$ breaks down and the curve must be reparametrized. For the metric (23), one finds
$\Delta=-\,\frac{4c_{d}}{d-2}\,r_{h}^{\,4-d}\neq 0.$ (99)
Appendix D Proof of the Positivity of $\partial_{P}\Phi_{1}$
$\partial_{P}\Phi_{1}=\frac{16\pi\,r_{\mathrm{isco}}}{(d-2)(d-1)}$ [ $\Bigl[8+4(d-1)(d-2)q^{2}r_{\mathrm{isco}}^{6-2d}$ (100)
$-c_{d}(d-1)(d+1)M\,r_{\mathrm{isco}}^{3-d}\Bigr].$
Since the prefactor is positive, the sign of $\partial_{P}\Phi_{1}$ is determined by the bracketed term. Let $x:=r_{\mathrm{isco}}^{\,d-3}>0$, so that
$B=8+\frac{4(d-1)(d-2)q^{2}}{x^{2}}-\frac{(d-1)(d+1)c_{d}M}{x}.$ (101)
Multiplying both sides by $x^{2}>0$ yields an upward-opening quadratic,
$\mathcal{Q}(x)=8x^{2}-(d-1)(d+1)c_{d}Mx+4(d-1)(d-2)q^{2},$ (102)
with discriminant
$\Delta_{x}=\big[(d-1)(d+1)c_{d}M\big]^{2}-128(d-1)(d-2)q^{2},$ (103)
and roots
$x_{\pm}=\frac{(d-1)(d+1)c_{d}M\pm\sqrt{\Delta_{x}}}{16},\qquad(\Delta_{x}\geq 0).$ (104)
Hence,
$\mathrm{sgn}\,\partial_{P}\Phi_{1}=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (105)
To determine the sign of $\partial_{P}\Phi_{1}$, we first examine the sign of $\Delta_{x}$. Under the isothermal constraint one has
$T_{0}=\frac{(d-1)c_{d}M\,r_{h}^{d+3}-2(d-2)q^{2}r_{h}^{6}-2r_{h}^{2d}}{4\pi r_{h}^{1+2d}}.$ (106)
The requirement $T_{0}>0$ is equivalent to
$(d-1)c_{d}M\,r_{h}^{d+3}-2(d-2)q^{2}r_{h}^{6}-2r_{h}^{2d}>0.$ (107)
Introducing $r_{h}^{2d}=r_{h}^{6}s^{2}$ and $r_{h}^{d+3}=r_{h}^{6}s$, we define
$\mathcal{B}(s):=-2s^{2}+(d-1)c_{d}\,M\,s-2(d-2)q^{2}>0,$ (108)
which is a downward-opening quadratic function. For $\mathcal{B}(s)>0$ to hold, its discriminant must satisfy
$\Delta_{s}=\big(c_{d}(d-1)M\big)^{2}-16(d-2)q^{2}>0,$ (109)
implying
$M>\frac{4\sqrt{d-2}}{c_{d}(d-1)}\,q.$ (110)
Substituting this inequality into Eq. (103) gives
$\Delta_{x}>0.$ (111)
Hence, $\partial_{P}\Phi_{1}$ may take positive, negative, or zero values, depending on the relative magnitude of $x$ and $x_{\pm}$.
To clarify the relation between $x$ and $x_{\pm}$, we rewrite
$\Phi_{1}(r,M,q,P,d)=P\,\partial_{P}\Phi_{1}+\widetilde{\Phi}_{1}(r,M,q,d).$ (112)
Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
$\partial_{P}\Phi_{1}\to 0.$ (113)
This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
$\Phi_{1}=\frac{2Mr_{\mathrm{isco}}^{2}(r_{\mathrm{isco}}-6M)+18Mq^{2}r_{\mathrm{isco}}-8q^{4}}{r_{\mathrm{isco}}^{5}}.$ (114)
The analysis can be divided into two cases:
  • $q=0$: For the Schwarzschild black hole, this yields $r_{\mathrm{isco}}=6M$. In this case, the two roots are $x_{\pm}={15M}/{4}$ and 0, and clearly $x\!\mid_{P=0}=6M>x_{+}$.
  • $q\neq 0$: For the RN black hole, it is well known that $r_{\mathrm{isco}}$ attains its minimum at $M=q$, where $r_{\mathrm{isco,min}}=4M$. From Eq. (104), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
Therefore, for both the Schwarzschild and RN black holes one always has $x\!\mid_{P=0}>x_{+}$, and consequently $\partial_{P}\Phi_{1}>0$ in this limit.
Next, consider increasing $P$ from zero in Eq. (112). Initially, one has $\Phi_{1}=\widetilde{\Phi}_{1}(r_{\mathrm{isco}}\!\mid_{P=0},M,q,d)=0$ with $\Gamma_{1}=\partial_{r}\widetilde{\Phi}_{1}>0$. Once $P$ increases, the term $\partial_{P}\Phi_{1}>0$ appears, so $\widetilde{\Phi}_{1}$ must decrease to maintain $\Phi_{1}=0$. Since $\Gamma_{1}>0$, this requires $r_{\mathrm{isco}}$ to decrease, meaning that $x$ moves toward $x_{+}$. As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to\infty}$.
Consequently, for any finite $P$, the corresponding $x$ must lie within the interval $(x_{+},\,x\!\mid_{P=0})$, which guarantees $\partial_{P}\Phi_{1}>0$ throughout. From a physical viewpoint, this reflects the suppressive effect of pressure on the ISCO radius: as $P$ increases, $r_{\mathrm{isco}}$ decreases monotonically toward $x_{+}$.
For higher dimensions ($d>4$), the argument becomes even simpler. At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to\infty$). As $P$ increases, $r_{\mathrm{isco}}$ moves inward from infinity toward $x_{+}$, again ensuring $\partial_{P}\Phi_{1}>0$. Hence, in all cases we conclude that
$\partial_{P}\Phi_{1}>0.$ (115)
Appendix E Asymptotic Relations in the Small-Horizon Limit
We now examine the asymptotic behavior of $V\,\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ in the small-horizon limit ($r_{h}\to 0^{+}$), focusing on their leading divergent terms.
For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix D for details). The thermodynamic volume takes the simple form
$V=\frac{\Omega_{d-2}\,r_{h}^{\,d-1}}{d-1},$ (116)
and therefore vanishes as $V\sim r_{h}^{\,d-1}\!\to\!0$. $\partial_{M}\Phi_{1}$ reads
$\partial_{M}\Phi_{1}=3c_{d}(d-1)(d-3)q^{2}r_{\mathrm{isco}}^{8-3d}\\ -2c_{d}^{2}(d-1)(d-3)M\,r_{\mathrm{isco}}^{5-2d}\\ -c_{d}(d-3)(d-5)r_{\mathrm{isco}}^{2-d}\\ -\frac{16c_{d}(d+1)\pi P\,r_{\mathrm{isco}}^{4-d}}{d-2},$ (117)
and its asymptotic behavior can be inferred once the leading scalings of $M$, $P$, and $r_{\mathrm{isco}}$ are determined.
Under the isothermal constraint, one finds
$M(r_{h})=\frac{\bigl(1+2\pi T_{0}r_{h}\bigr)\,r_{h}^{2(d-3)}+2(d-2)\,q^{2}}{c_{d}\,(d-1)\,r_{h}^{\,d-3}},$ (118)
and
$P(r_{h})=\frac{(d-2)\!\left[\bigl(3-d+4\pi T_{0}r_{h}\bigr)\,r_{h}^{2(d-3)}+(d-3)\,q^{2}\right]}{16\pi\,r_{h}^{\,2(d-2)}}.$ (119)
In the small-horizon limit ($r_{h}\!\to\!0^{+}$), these behave as
$M\sim\frac{1}{r_{h}^{\,d-3}},\qquad P\sim\frac{1}{r_{h}^{\,2(d-2)}}.$ (120)
For the ISCO, the divergence $P\!\to\!\infty$ implies (see Appendix D)
$r_{\mathrm{isco}}^{\,d-3}\to x_{+}=\frac{(d-1)(d+1)c_{d}M+\sqrt{\Delta_{x}}}{16},$ (121)
where
$\Delta_{x}=\big[(d-1)(d+1)c_{d}M\big]^{2}-128(d-1)(d-2)q^{2}.$ (122)
Near $r_{h}\!\to\!0^{+}$, the dominant contribution arises from the mass term $M$, so that
$r_{\mathrm{isco}}\sim M^{\tfrac{1}{d-3}}\sim\frac{1}{r_{h}}.$ (123)
Substituting Eqs. (120) and (123) into Eq. (117) gives $\partial_{M}\Phi_{1}\sim-r_{h}^{-d}$, and hence
$V\,\partial_{M}\Phi_{1}\sim-\frac{1}{r_{h}}\;\longrightarrow\;-\infty.$ (124)
Appendix F Monotonicity of $r_{\rm c}(r_{h})$ in the Phase-Transition Region
F.1 Photon sphere ($\epsilon=0$)
Our proof consists of two steps. First, we demonstrate that $r_{h0}$ lies to the left of $r_{h,\min}$. Second, we show that $P(r_{h0})>P(r_{h,\max})$, where $r_{h,\min}$ and $r_{h,\max}$ denote the locations of the minimum and maximum of $P(r_{h})$, respectively. This procedure circumvents the need to determine the small-black-hole radius $r_{h,s}$ from Maxwell’s equal-area law while still guaranteeing that $r_{h0}$ lies to the left of $r_{h,s}$.
Step 1
From Eq. (41) it follows that ${\mathrm{d}P}/{\mathrm{d}r_{h}}=0$ implies ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$. Since $M(r_{h})$ first decreases and then increases, $r_{h0}$ must therefore lie to the left of both extrema, namely on the small-black-hole branch.
Step 2
To establish the result for arbitrary $q$, we eliminate $q$ by expressing $P(r_{h})$ and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ in reduced variables:
$P(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (125)
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (126)
where $\tilde{r}_{h}=r_{h}/r_{h,c}$ and $\tilde{T}_{0}=T_{0}/T_{c}$, with $r_{h,c}=\sqrt{6}\,q$ and $T_{c}=1/(3\sqrt{6}\,\pi q)$ denoting the critical values. Using these relations, we numerically determine $r_{h0}$ and $r_{h,\max}$, and plot $P(r_{h0})$ (red curve) and $P(r_{h,\max})$ (blue curve) as functions of $\tilde{T}$ in Fig. 9. The black dashed vertical line at $\tilde{T}_{0}=\sqrt{2}/2$ marks the critical value above which $P>0$; the region $\tilde{T}_{0}<\sqrt{2}/2$ is therefore unphysical and excluded from consideration. From the remaining portion of the plot it is evident that $P(r_{h0})>P(r_{h,\max})$. Combined with the result of Step 1, this demonstrates that $r_{h0}$ lies to the left of $r_{h,s}$.
To visualize this behavior, Fig. 10 shows $P(r_{h})$ (red) and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ (blue) for several $\tilde{T}_{0}$. At $\tilde{T}_{0}=0.4$, $P(r_{h0})<P(r_{h,\max})$, but $P(r_{h})<0$ and is thus unphysical. When $\tilde{T}_{0}\geq\sqrt{2}/2$, $P(r_{h})\geq 0$ and $P(r_{h0})>P(r_{h,\max})$, with $r_{h,s}$ and $r_{h,l}$ clearly located to the right of $r_{h0}$.
F.2 ISCO ($\epsilon=1$)
With the photon-sphere case established, the ISCO analysis becomes straightforward. From the criterion
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right),$ (127)
we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is momentarily ignored, the zeros of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ coincide with those of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{0}$, since both are controlled entirely by $\mathrm{d}M/\mathrm{d}r_{h}$. Because $\partial_{M}\Phi_{1}<0$ and $\mathrm{d}M/\mathrm{d}r_{h}$ increases monotonically, ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ is strictly decreasing.
Restoring the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}<0$ can only shift the zero of $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}$further to the left. Since the zero of $\mathrm{d}M/\mathrm{d}r_{h}$ already lies to the left of $r_{h,s}$ (as shown earlier), its additional leftward shift necessarily keeps it on the small-black-hole side.
Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN-AdS black hole is globally non-monotonic, it remains locally monotonic within the phase-transition region, thereby preserving the discontinuous jump across the transition.
Figure 9: Plots of $p(r_{h0})$ (red) and $p(r_{h,\max})$ (blue) as functions of $\tilde{T}_{0}$. At $\tilde{T}_{0}=\sqrt{2}/2$, one has $p(r_{h})>0$ for all $r_{h}$.
Figure 10: Behaviors of $P(r_{h})$ and $\mathrm{d}M/\mathrm{d}r_{h}$ for three representative values of $\tilde{T}_{0}$.

### v2

@@
 Abstract
-In recent years, jumps in the particle circular-orbit radius $r_{\rm c}$ have been proposed as observational signatures of black hole phase transitions, supported by the general consensus that $r_{\rm c}$ increases monotonically with the horizon radius $r_{h}$.
-However, in this paper, we will show that this monotonicity is not universally valid.
-By deriving a universal, ensemble-independent criterion for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ and applying it to the $d$-dimensional Reissner-Nordström-Anti-de Sitter black hole, we reveal a sharp dichotomy: while $r_{\rm c}(r_{h})$ exhibits strictly monotonicity along isobars, it typically becomes non-monotonic along isotherms (except for the photon sphere of Schwarzschild-Anti-de Sitter black hole)–a behavior that has been overlooked in previous studies.
-Nevertheless, we rigorously prove that within the phase-transition region itself, the monotonicity of $r_{\rm c}(r_{h})$ is preserved.
-However, the generalizability of this result to other spacetimes requires case-by-case verification, and its validity could have notable implications for the observability of the jump in $r_{\rm c}$.
-Interestingly, non-monotonicity itself acts as a new observational probe: the monotonicity contrast in $r_{\rm c}(r_{h})$ distinguishes isobaric from isothermal evolution, providing a complementary thermodynamic probe.
-Furthermore, we show that the critical exponents extracted from $r_{\rm c}$ match those obtained from $r_{h}$, independent of ensemble.
-Remarkably, all these phenomena are closely tied to the thermodynamic first law.
+Recent studies propose that black hole phase transitions can be encoded in the circular orbit radius of particles.
+In this paper, we systematically investigate the reliability of this encoding mechanism.
+We find that this mechanism is highly reliable in the isobaric ensemble, whereas it may break down in the isothermal ensemble.
+It turns out that the reliability of this mechanism is directly controlled by the first law of black hole thermodynamics.
+Interestingly, even if this encoding mechanism fails, we prove that, for any black hole exhibiting criticality, the first law can ensure that, near the critical point, the coexistence gap of the circular orbit radius remains a reliable order parameter and yields exactly the same critical exponents as the standard thermodynamic order parameter.
+Our results provide a potential way to identify the thermodynamic ensemble of a black hole, and reveal a deeper connection between gravitational geometry and thermodynamics.
 I Introduction
-Black hole thermodynamics has evolved into a central theme of gravitational physics.
-The Hawking-Page transition first revealed a first-order transition between thermal Anti-de Sitter (AdS) radiation and large AdS black holes [1].
-Subsequent work has uncovered thermodynamic phase transitions in various black-hole spacetimes, including Van der Waals-type small/large transitions in AdS black holes [2, 3, 4, 5, 6, 7], thereby establishing a rich thermodynamic phase structure.
-Yet direct observation of these thermodynamic phenomena remains limited.
-In contrast, optical and dynamical features–such as the photon sphere and the innermost stable circular orbit (ISCO)–are directly observable and encode essential geometric information relevant to shadows and accretion physics [8, 9, 10].
-The EHT image of M87* made this explicit, revealing an emission ring whose boundary traces the photon orbit [11].
-This naturally raises the possibility that black hole thermodynamics could be probed through optical features.
-Compelling evidence comes from the work of Wei et al., who showed that in charged AdS black holes, both the photon-sphere radius and the minimum impact parameter exhibit Van derWaals-type non-monotonicity and jump discontinuities, behaving as mean-field order parameters [12].
-Subsequently, related signatures involving photon orbits and ISCOs have been identified across a variety of spacetimes and modified gravitational theories [13, 14, 15].
-A central consideration in these studies is the relationship between the circular-orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$.
-Across a first-order small/large transition, $r_{h}$ undergoes a discontinuous jumps, and the monotonicity $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly–making it a natural geometric order parameter (see Appendix A for a brief mathematical derivation).
-This expectation was later formalized by Zhang et al., who proved monotonic growth of both radii (the photon-sphere radius and ISCO radius) with $r_{h}$ in general static spherical geometries–a result that has been widely applied in subsequent work [16].
-However, we find that once more than one thermodynamic parameter is allowed to vary, this monotonicity claim fails.
-To illustrate this, we present a simple counterexample based on the Schwarzschild-AdS (Sch-AdS) black hole: the photon-sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and $P$, where $P=-\Lambda/(8\pi)$ with $\Lambda$ being the negative cosmological constant.
+Black hole thermodynamics provides a unique arena for exploring the interplay among gravity, quantum physics, and statistical mechanics.
+It began as a formal analogy between the laws of black hole mechanics and ordinary thermodynamics, and acquired genuine physical content through the generalized second law and Hawking radiation [1, 2, 3, 4].
+Subsequent developments connected it to broader areas of physics, for example microscopic state counting in string theory and holography in AdS/CFT [5, 6, 7].
+Particularly influential progress has come from AdS black holes in the extended phase space, where the cosmological constant is interpreted as a pressure and the mass as an enthalpy.
+Within this framework, AdS black holes were found to possess a rich phase structure, including Van der Waals-like phase transitions and criticality, as well as more exotic behaviors in modified gravity and nonlinear electrodynamics [8, 9, 10, 11, 12, 13, 14, 15, 16].
+These advances further confirm that black hole thermodynamics carries genuine physical content and naturally raise the question how it can be tested.
+A particularly compelling possibility comes from horizon-scale imaging.
+The Event Horizon Telescope has imaged the shadows of the supermassive black holes in M87 and at the Galactic center, opening a new observational window on strong-field optical phenomena [17, 18].
+Motivated by this breakthrough, recent theoretical work has begun to explore how thermodynamic phase transitions may leave imprints on optical observables.
+In particular, Wei et al.
+showed that in the isobaric ensemble the photon sphere radius can serve as an effective probe of the small-/large-black hole phase transition, suggesting that thermodynamic evolution may be reflected in observable image features [19, 20].
+Subsequently, related signatures involving photon spheres and innermost stable circular orbits (ISCOs) have been identified across a variety of spacetimes and modified theories of gravity [21, 22, 23, 24].
+A key ingredient in these studies is the monotonicity of the mapping between the circular orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$.
+Across a first-order small/large phase transition, $r_{h}$ undergoes a discontinuous jump, and the monotonic mapping $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly, thereby making $r_{\rm c}$ a natural geometric order parameter (see Appendix A for a brief mathematical derivation).
+This expectation was later formalized by Zhang et al., who proved a monotonic increase of the photon sphere and ISCO radii with $r_{h}$ in general static, spherically symmetric geometries, and this result has been widely adopted in subsequent work [25, 26].
+However, we find that once more than one thermodynamic parameter is allowed to vary, the claim of monotonic increase may fail.
+To illustrate this, we present a simple counterexample based on the Schwarzschild–AdS (Sch–AdS) black hole: the photon sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and the pressure $P=-\Lambda/(8\pi)$ associated with the negative cosmological constant $\Lambda$.
 Increasing $P$ decreases $r_{h}$, with $r_{\rm ps}$ unchanged.
 Consequently, there exist smooth paths in $(M,P)$ parameter space along which $r_{h}$ increases and $r_{\rm ps}$ decreases.
-Figure 1 provides an explicit example, and Appendix B presents a fully analytic realization.
-This breakdown of monotonicity in $r_{\rm c}(r_{h})$ has been neglected in previous studies.
-The reason is straightforward: most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically grow monotonically with $M$.
-Even in studies that employed the isothermal ensembles (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not systematically examined, leading to the omission of non-monotonic behavior.
-Thus, a genuine understanding of how circular orbits encode thermodynamic information requires analyzing the monotonicity of $r_{\rm c}(r_{h})$ along arbitrary parameter paths–precisely the general framework developed in this work.
-Figure 1: Behavior of $M(r_{h},P)$ in the Sch-AdS spacetime.
-Thin black solid curves denote contours of constant mass $M$.
-Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon-sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
+Figure 1 provides an explicit path example, and Appendix B presents an analytic example by constructing an explicit parametric path.
+This breakdown of monotonic increase in $r_{\rm c}(r_{h})$ has been neglected in previous studies, and the reason is straightforward: Most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically increasing monotonically with $M$.
+Even when isothermal ensembles were considered (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not examined systematically, so possible monotonic decrease or nonmonotonicity was overlooked.
+Thus, understanding how circular orbits encode black hole phase transitions and other thermodynamic information calls for analyzing the monotonicity of $r_{\rm c}(r_{h})$ along general parameter paths, which is precisely the focus of this work.
 Our strategy is as follows.
 In Sec.
-II, we review the conditions for photon spheres and ISCOs in a general static, spherically symmetric spacetime, and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary parameter path.
+II, we review the conditions for the photon sphere and the ISCO in a general static, spherically symmetric spacetime and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary path in parameter space.
 In Sec.
-III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström-AdS (RN-Ads) black holes, focusing on the isobaric ($T$-$S$) and isothermal ($P$-$V$) thermodynamic paths.
+III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström–AdS (RN–AdS) black holes, focusing on isobaric and isothermal thermodynamic paths.
 In Sec.
-IV, we show that the critical exponents remain invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, demonstrating that $r_{\rm c}$ defines the same universality class as $r_{h}$.
+IV, we show that the critical exponents are invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, so that $r_{\rm c}$ and $r_{h}$ yield the same universality class.
 Finally, Sec.
 V summarizes and discusses our main results.
+Throughout this work we employ natural units with $c=G=1$.
+Figure 1: Behavior of $M(r_{h},P)$ in the Sch–AdS spacetime.
+Thin black solid curves denote contours of constant mass $M$.
+Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
 II Monotonicity criterion
@@
 $\mathrm{d}s^{2}=-f(r;\lambda)\,\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{f(r;\lambda)}+r^{2}\mathrm{d}\Omega^{2},$ (1)
-where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black-hole parameters, such as mass $M$, charge $q$, and so on.
+where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black hole parameters, such as mass $M$, charge $q$, and so on.
 The Lagrangian of a particle moving in the spacetime determined by Eq.
@@
 The particle’s circular orbits at $r=r_{c}$ are determined by
-$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0.$ (6)
-Solving Eq. (6) for a massless particle with $\epsilon=0$ gives
+$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0,$ (6)
+where the prime denotes differentiation with respect to $r$.
+Solving Eq. (6) for a massless particle with $\epsilon=0$ gives (see Appendix C for a detailed derivation)
 $\Phi_{0}(r;\lambda):=r\,\partial_{r}f(r;\lambda)-2f(r;\lambda)=0,\quad\text{at}\,r=r_{\rm ps},$ (7)
-where $r_{\rm ps}$ denotes the photon-sphere radius. The condition for an unstable photon sphere further requires
-$\mathcal{R}^{\prime\prime}(r_{\rm ps})<0,$ (8)
+where $r_{\rm ps}$ denotes the photon sphere radius. The condition for an unstable photon sphere further requires
+$\mathcal{R}^{\prime\prime}(r_{\rm ps})>0,$ (8)
 which is equivalent to
@@
 and the corresponding stability criteria
-$\Gamma_{\epsilon}(\lambda)\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
+$\Gamma_{\epsilon}\,\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
 II.2 Derivation of the Monotonicity Criterion
-To investigate the monotonic relation between the circular-orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
+To investigate the monotonic relation between the circular orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
 $\Phi_{\epsilon}(r_{c};\lambda)=0,\qquad f(r_{h};\lambda)=0.$ (15)
@@
 and
-$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa>0,$ (18)
+$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa\,\neq\,0,$ (18)
 where $\kappa=\tfrac{1}{2}\,\partial_{r}f(r;\lambda)\big|_{(r_{h};\lambda)}$ is the surface gravity of the black hole, which is positive for any nonextremal case.
@@
 (15) can in principle be differentiated with respect to $\lambda$.
-Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(t)$ in the parameter space, along which all quantities become functions of a single real variable $t$.
+Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(\tau)$ in the parameter space, along which all quantities become functions of a single real variable $\tau$.
 This allows us to directly apply the chain rule, yielding
-$\frac{\mathrm{d}}{\mathrm{d}t}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(t);\lambda(t)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
-$\frac{\mathrm{d}}{\mathrm{d}t}\,f\big(r_{h}(t);\lambda(t)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
-where the overdot denotes differentiation with respect to $t$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$.
+$\frac{\mathrm{d}}{\mathrm{d}\tau}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(\tau);\lambda(\tau)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
+$\frac{\mathrm{d}}{\mathrm{d}\tau}\,f\big(r_{h}(\tau);\lambda(\tau)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
+where the overdot denotes differentiation with respect to $\tau$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$.
 Combining Eqs.
@@
 It therefore determines whether the overall monotonicity of $r_{\mathrm{c}}(r_{h})$ is preserved or violated, emphasizing the importance of considering multiparameter constraints in black hole thermodynamics, as discussed in the Introduction.
-In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN-AdS black hole.
+In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN–AdS black hole.
 III Examples of $d$-Dimensional Black Holes under Isobaric and Isothermal Constraints
@@
 depend only on the spacetime dimension, with $\Omega_{d-2}$ being the area of the unit (d2)(d-2)-sphere. The metric (23) reproduces several well-known limits:
-  • $q=0$: the $d$-dimensional Sch-AdS black hole;
-  • $q\neq 0$: the RN-AdS black hole;
+  • $q=0$: the $d$-dimensional Sch–AdS black hole;
+  • $q\neq 0$: the RN–AdS black hole;
   • $P\to 0$: the asymptotically flat Schwarzschild or RN solution in $d$ dimensions.
@@
 This leads to $\dot{r}_{h}={\mathrm{d}r_{h}}/{\mathrm{d}M}$.
-To examine whether $\dot{r}_{h}$ can vanish, we consider the thermodynamic first law
+To examine whether $\dot{r}_{h}$ can vanish, we consider the first law of black hole thermodynamics
 $\mathrm{d}M=T\,\mathrm{d}S\quad\Rightarrow\quad\frac{\mathrm{d}r_{h}}{\mathrm{d}M}=\frac{1}{T\,S^{\prime}(r_{h})},$ (25)
@@
 The monotonicity of $r_{\rm c}(r_{h})$ is therefore governed by $\partial_{M}\Phi_{\epsilon}$.
-Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ determines the background geometry and hence controls the response of circular orbits, which generically vary with $M$.
+Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ fixes the background geometry and thus controls the circular orbits, so they generically vary with $M$.
+Based on this consideration and the universality of the first law, we expect that for almost all black holes the mapping $r_{\rm c}(r_{h})$ is monotonic in the isobaric ensemble.
+As a result, when a phase transition drives the horizon radius to jump from the small-black hole branch $r_{h,s}$ to the large-black hole branch $r_{h,l}$, the circular orbit radius undergoes a corresponding jump from $r_{\rm c}(r_{h,s})$ to $r_{\rm c}(r_{h,l})$.
+This makes the circular orbit radius a robust and reliable encoder of the phase transition information.
 For the metric (23), we obtain
@@
 $\partial_{M}\Phi_{1}$ $=-\frac{2\,r_{\rm isco}^{\,1-4d}}{(d-1)(d-2)}\,\sqrt{h(r_{\rm isco},q,P,d)}<0,$
-where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into Eq. (27) and using Eqs. (14) and (26), we find
+where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into the criterion (27) and using Eqs. (14) and (26), we find
 $\epsilon=0:$ $\quad-\frac{1}{\Gamma_{0}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}>0,\ \partial_{M}\Phi_{0}>0\ \Rightarrow\ \frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}>0,$ (29)
@@
 The opposite signs of $\Gamma_{0}$ and $\Gamma_{1}$ are exactly compensated by those of $\partial_{M}\Phi_{0}$ and $\partial_{M}\Phi_{1}$, resulting in a consistent monotonic relation between $r_{\rm c}$ and $r_{h}$.
-This implies that, during an isobaric process, a small-to-large black hole phase transition necessarily induces a corresponding discontinuous jump in both $r_{\mathrm{ps}}$ and $r_{\mathrm{isco}}$.
-Such abrupt changes could, in principle, be detected through astronomical observations, providing a potential signature of black hole phase transitions.
 These findings are consistent with the earlier results reported by Wei, Zhang, et al.
-Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional (4D) RN-AdS black hole, one obtains
+Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional RN–AdS black hole, one obtains
 $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{2\kappa}{{c_{d}}/{r_{h}^{\,d-3}}}=\frac{2(-{q^{2}}/{r_{h}^{3}}+{1}/{r_{h}}+8\pi P\,r_{h})}{{c_{d}}/{r_{h}^{\,d-3}}}>0,$
@@
 […]
-It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq. (25). Using Eq. (27) together with the above results, we find
+It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq.
+(25).
+Using the criterion (27) together with the above results, we find
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3\,r_{\rm ps}^{2}\!\left(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4}\right)}{2\left(2q^{2}-r_{\rm ps}^{2}\right)r_{h}^{2}},$ (33)
@@
 $\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=-\,6\,(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4})\,r_{\rm isco}^{13}\,\sqrt{g}\;/\\ \Bigl[r_{h}^{2}\bigl(15q^{4}r_{\rm isco}^{11}+r_{\rm isco}^{15}+352\pi P\,r_{\rm isco}^{17}+4800\pi^{2}P^{2}r_{\rm isco}^{19}\\ +48\pi P\,q^{2}\,r_{\rm isco}^{15}+(r_{\rm isco}^{2}-120\pi P\,r_{\rm isco}^{4}-9q^{2})\,r_{\rm isco}^{11}\,\sqrt{g}\bigr)\Bigr].$ (34)
-In the Sch-AdS limit ($q=0$), the above relations reduce to
+In the Sch–AdS limit ($q=0$), the above relations reduce to
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2}+12\pi Pr_{h}^{2},$ (35)
@@
 […]
-where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further taking the Schwarzschild limit ($q=0$, $P=0$), one obtains
+where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further, in the limit where both $q$ and $P$ approach zero, one obtains
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2},\qquad\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=3,$ (37)
-which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$.
+which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$ in the Schwarzschild case.
 In Figs.
 2 and 3, we present the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$ and $r_{\mathrm{isco}}(r_{h})$ for the four-dimensional case.
-As shown, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch-AdS black hole, whereas for the RN-AdS black hole it first decreases and then increases, with the extremum occurring precisely at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$.
+In both panels, $P_{c}$ denotes the pressure at the RN–AdS critical point, obtained from the criticality conditions $\partial T/\partial r_{h}=\partial^{2}T/\partial r_{h}^{2}=0$.
+As illustrated, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch–AdS black hole, whereas for the RN–AdS black hole, it first decreases and then increases, with the extremum exactly at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$.
 Truncating the curve at this point, $r_{\mathrm{c}}(r_{h})$ retains its local monotonicity.
@@
 Figure 2: Behavior of $r_{\rm ps}(r_{h})$ along isobaric paths in $d=4$.
-For the Sch-AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$ corresponding to the $q=1$ reference point.
-The same choice is adopted for subsequent figures and will not be repeated.
+For the Sch–AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$, where $P_{\mathrm{c}}$ is the critical pressure of the RN–AdS solution with $q=1$.
+The same choice is adopted for subsequent figures and will not be restated hereafter.
 Figure 3: Behavior of $r_{\rm isco}(r_{h})$ along isobaric paths in $d=4$
@@
 The isothermal process corresponds to the $P$-$V$ ensemble commonly employed in black hole thermodynamics.
-Here the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$.
+Here, the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$.
 Therefore, we take $\lambda=(M,P;\lambda_{\mathrm{fixed}})$.
-Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter gives (see Appendix C for justification)
+Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter yields the following result [see Appendix D for justification of the existences of $M(r_{h})$ and $P(r_{h})$]
 $\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right).$ (38)
 In general, this relation has no simple closed form.
-However, for photons ($\epsilon=0$), many black-hole metrics satisfy $\partial_{P}\Phi_{0}=0$, for example when the pressure enters the metric only through a term proportional to $Pr^{2}$.
+However, for photons ($\epsilon=0$), many black hole metrics satisfy $\partial_{P}\Phi_{0}=0$, e.g., when the pressure enters the metric only through a term proportional to $Pr^{2}$.
 In such cases, the criterion (38) reduces to
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{0}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\partial_{M}\Phi_{0}.$ (39)
-This result is consistent with the criterion for the isobaric process, but differs in that under the isobaric constraint the first law of thermodynamics requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}\neq 0$, ensuring that $r_{\mathrm{ps}}(r_{h})$ remains monotonic.
-In contrast, under the isothermal constraint the first law takes the form
+This result is consistent with the criterion for the isobaric process, but differs in that the isobaric constraint requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$ via the first law, ensuring the monotonicity of $r_{\mathrm{ps}}(r_{h})$.
+In contrast, under the isothermal constraint, the first law takes the form
 $\mathrm{d}M=T_{0}\,\mathrm{d}S+V\,\mathrm{d}P,$ (40)
@@
 $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (41)
-indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become non-monotonic.
+indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish or even become negative, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become nonmonotonic.
 This behavior is physically intuitive: when ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$, one must have
@@
 This means that, near an extremum of $M(r_{h})$, the same $M$ corresponds to two distinct $P$.
-As $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, consequently a single $r_{\rm ps}$ maps to two different $r_{h}$, and $r_{\rm ps}(r_{h})$ is necessarily non-monotonic.
-Next, we apply the criterion (38) to the metric (23).
+Since $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, a single $r_{\rm ps}$ maps to two different $r_{h}$, and thus $r_{\rm ps}(r_{h})$ is necessarily nonmonotonic.
+For massive particles ($\epsilon=1$), one typically has $\partial_{P}\Phi_{1}\neq 0$.
+Taking into account that $M$ and $P$ usually affect $r_{\rm isco}$ in opposite ways, we assume that $\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ carry opposite signs (as is indeed the case for RN–AdS black holes).
+Under this assumption, the criterion (38) shows that if $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ have opposite signs, then $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ cannot cross zero, so $r_{\rm isco}(r_{h})$ remains monotonic.
+Conversely, if $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ share the same sign, $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ can in principle vanish, leading to a nonmonotonic $r_{\rm isco}(r_{h})$.
+Interestingly, the first law naturally allows for such nonmonotonic behavior: from Eq.
+(41) one sees that $\mathrm{d}P/\mathrm{d}r_{h}>0$ implies $\mathrm{d}M/\mathrm{d}r_{h}>0$, while $\mathrm{d}M/\mathrm{d}r_{h}<0$ implies $\mathrm{d}P/\mathrm{d}r_{h}<0$, so there exist regimes where $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ have the same sign and $r_{\rm isco}(r_{h})$ can become nonmonotonic.
+Conversely, if we assume that $\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ share the same sign, the first law will require that $r_{\rm isco}(r_{h})$ be nonmonotonic.
+The reason is simple: At an extremum of $M(r_{h})$, one has $\mathrm{d}M/\mathrm{d}r_{h}=0$, and thus Eq.
+(41) requires $\mathrm{d}P/\mathrm{d}r_{h}<0$.
+Substituting these relations into the criterion (38) yields
+$\left.\mathrm{sgn}\!\left(\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}\right)\right|_{\mathrm{d}M/\mathrm{d}r_{h}=0}=\mathrm{sgn}\!\left(-\frac{1}{\Gamma_{1}}\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}\right).$ (43)
+At an extremum of $P(r_{h})$, one instead has $\mathrm{d}P/\mathrm{d}r_{h}=0$, and thus Eq. (41) gives $\mathrm{d}M/\mathrm{d}r_{h}>0$. Inserting this into the criterion (38) leads to
+$\left.\mathrm{sgn}\!\left(\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}\right)\right|_{\mathrm{d}P/\mathrm{d}r_{h}=0}=\mathrm{sgn}\!\left(-\frac{1}{\Gamma_{1}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}\right).$ (44)
+It is easy to see that the two above expressions have different signatures.
+Therefore, the first law enforces a sign change in $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$, implying that $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic, with its extremum located between those of $M(r_{h})$ and $P(r_{h})$.
+Whether such a configuration is actually realized depends on the specific black hole model and must be examined case by case.
+In any case, these analyses show that, in the isothermal ensemble, the first law no longer enforces $r_{c}(r_{h})$ to be monotonic, but instead allows nonmonotonic behavior.
+Once such nonmonotonicity arises, the jump $\Delta_{c}=r_{c}(r_{h,l})-r_{c}(r_{h,s})$ can be strongly suppressed or even vanish, thereby undermining the reliability of circular orbit radii as encoders of phase transitions.
+Next, we apply criterion (38) to the metric (23) to verify the above analysis explicitly.
 III.2.1 Photon sphere ($\epsilon=0$)
@@
 so ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is strictly increasing in $r_{h}$. Meanwhile, From Eq. (45),
-$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (46)
-hence there exists a unique $r_{h0}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h0}}=0$. By the criterion (39), it follows that:
+$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to+\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (48)
+hence there exists a unique $r_{h,e_{1}}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h,e_{1}}}=0$. By the criterion (39), it follows that:
   • $q=0$: $r_{\rm ps}(r_{h})$ is monotonically increasing;
-  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is non-monotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h0}$, and then increases.
-This result reveals that the presence of electric charge $q$ breaks the monotonic behavior of $r_{\mathrm{ps}}(r_{h})$, making it globally non-monotonic–an aspect that has been omitted in previous studies.
+  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is nonmonotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h,e_{1}}$, and then increases.
+These results confirm our previous analysis that $r_{\rm ps}(r_{h})$ can indeed become nonmonotonic in the isothermal ensemble, which has not been noticed in earlier studies.
+Interestingly, in the Sch–AdS case ($q=0$), $r_{\rm ps}(r_{h})$ remains monotonic.
+This can be readily understood from Eq.
+(41): When $q=0$, the slope $\mathrm{d}P/\mathrm{d}r_{h}$ is not sufficiently negative to drive $\mathrm{d}M/\mathrm{d}r_{h}$ across zero.
+By contrast, for the RN–AdS case ($q\neq 0$), $P(r_{h})$ exhibits a Van der Waals–like behavior, and at small $r_{h}$ its slope $\mathrm{d}P/\mathrm{d}r_{h}$ becomes strongly negative, which is sufficient to make $\mathrm{d}M/\mathrm{d}r_{h}$ cross zero.
 The explicit expression for ${dr_{\mathrm{ps}}}/{dr_{h}}$ reads
@@
 $\quad\Big[(d-3)\big((2-d)\,q^{2}\,r_{\rm ps}^{6}+r_{\rm ps}^{2d}\big)\,r_{h}^{\,4+d}\Big],$
-For $d=4$ (the RN-AdS case), this reduces to
+For $d=4$ (the RN–AdS case), this reduces to
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{r_{\rm ps}^{2}\,\left[2q^{2}-r_{h}^{2}\,(1+4\pi T_{0}\,r_{h})\right]}{(2q^{2}-r_{\rm ps}^{2})\,r_{h}^{2}},$ (50)
-and, in the Sch-AdS limit ($q=0$), further simplifies to
+and, in the Sch–AdS limit ($q=0$), further simplifies to
 $\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=1+4\pi T_{0}\,r_{h}.$ (51)
-This coincides with the standard Sch-AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
-Figure 4 presents the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$.
-For the Sch-AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$ and increases monotonically thereafter, remaining globally monotonic in full agreement with Eq.
+This coincides with the standard Sch–AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
+To verify the above analysis, we have performed numerical calculations and displayed the results in Fig.
+4.
+In this figure, $T_{c}$ denotes the temperature at the RN–AdS critical point, obtained from the criticality conditions $\partial P/\partial r_{h}=\partial^{2}P/\partial r_{h}^{2}=0$.
+For the Sch–AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$, where $r_{h,\min}$ is defined by $P(r_{h,\min})=0$, and smaller $r_{h}$ would give $P<0$ and are therefore discarded as unphysical.
+It then increases monotonically, remaining globally monotonic in full agreement with Eq.
 (51).
-In contrast, the RN-AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
-The breakdown of monotonicity naturally raises an important question: does it affect the discontinuous jump of $r_{\mathrm{ps}}$ across the phase transition, potentially obscuring its optical signature?
-The answer is: possibly.
-If the small and large black hole radii, $r_{h,s}$ and $r_{h,l}$, lie on opposite sides of the extremum of $r_{\mathrm{ps}}(r_{h})$, the amplitude of the discontinuous change in $r_{\mathrm{ps}}$ may be reduced, requiring higher observational resolution to detect it.
-In the extreme case where $r_{\mathrm{ps}}(r_{h,s})=r_{\mathrm{ps}}(r_{h,l})$, the discontinuity would vanish entirely, rendering the phase transition optically unobservable.
-From a thermodynamic perspective, however, this is not necessarily a drawback.
-As shown earlier, $r_{\mathrm{c}}(r_{h})$ remains monotonic in the isobaric process, while $r_{\mathrm{ps}}(r_{h})$ becomes non-monotonic in the isothermal one.
-Such a contrast in optical behavior between the two ensembles offers a potential diagnostic for identifying the underlying thermodynamic process of a black hole– a perspective that provides a new avenue for probing black hole thermodynamics, to be explored in future work.
-Another possibility is that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{ps}}(r_{h})$, so that $r_{\mathrm{ps}}(r_{h})$ remains locally monotonic throughout the phase-transition region.
-This scenario indeed occurs for the RN-AdS black hole studied here, and in Appendix F we provide an explicit proof for the $d=4$ case.
-Figure 5 shows the numerically computed profiles of $P(r_{h})$ and $P(r_{\mathrm{ps}})$ at fixed parameters, where the characteristic S-shaped behavior of the phase transition is fully preserved.
-In this configuration, one can observe both the global non-monotonicity of $r_{\mathrm{ps}}(r_{h})$ and the discontinuous jump of $r_{\mathrm{ps}}$ across the transition.
-The coexistence of these two optical signatures provides an especially powerful means for extracting thermodynamic information from black holes.
+In contrast, the RN–AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
 Figure 4: Behavior of $r_{\rm ps}(r_{h})$ along isothermal paths in $d=4$.
-The marked point on the curve corresponds to $P=0$ at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated.
-For the Sch-AdS case, the temperature is likewise set to the value $T=0.8\,T_{\mathrm{c}}$ corresponding to the $q=1$ reference point.
+The marked point on the curve corresponds to $P=0$ at $r_{h,\min}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated.
+For the Sch–AdS case, we likewise set the temperature to $T=0.8\,T_{\mathrm{c}}$, where $T_{\mathrm{c}}$ is the critical temperature of the RN–AdS solution with $q=1$.
 The same choice is adopted for subsequent figures and will not be repeated.
-Figure 5: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$.
 III.2.2 ISCO ($\epsilon=1$)
@@
 $-c_{d}(d-1)(d+1)M\,r_{\mathrm{isco}}^{3-d}\Bigr],$
-where $\partial_{P}\Phi_{1}>0$ (see Appendix D for proof).
-Since van der Waals-type behavior occurs only for charged black holes ($q\neq 0$), we analyze the RN-AdS and Sch-AdS ($q=0$) cases separately.
+where $\partial_{P}\Phi_{1}>0$ (see Appendix E for proof).
+Since van der Waals-type behavior occurs only for charged black holes, we analyze the RN–AdS and Sch–AdS cases separately.
 Moreover, because the explicit expression for ${\mathrm{d}r_{\mathrm{c}}}/{\mathrm{d}r_{h}}$ is algebraically cumbersome, we focus on whether it changes sign rather than displaying the full formula.
-(i) Sch-AdS ($q=0$)
+(i) Sch–AdS ($q=0$)
 For $q=0$, Eq. (52) reduces to
 $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\,\frac{(d-2)\,\bigl(3-d+2\pi T_{0}r_{h}\bigr)}{8\pi\,r_{h}^{3}}.$ (54)
-It has a single zero at $r_{h1}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h1}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h1}$.
+It has a single zero at $r_{h,e_{2}}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h,e_{2}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h,e_{2}}$.
 Using the criterion (38), together with ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}<0$, $\Gamma_{1}>0$, and $\partial_{P}\Phi_{1}>0$, one thus expects $r_{\rm isco}(r_{h})$ to first decrease and then increase.
@@
 For sufficiently large $r_{h}$ one has ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, which forces $\delta<0$.
-It remains to verify that $\delta>0$ for small $r_{h}$, ensuring a single zero crossing.
+It remains to verify that $\delta>0$ for small $r_{h}$, ensuring that $\delta$ crosses zero at least once.
 In this spirit we find
-$\begin{cases}\delta\bigl(3r_{h_{\min}},\,r_{h_{\min}},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h_{\min}},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (54)
-where $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h_{\min}})=0$.
-For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h_{\min}}$.
-For $d>4$ and $P\to 0$, the higher-dimensional Schwarzschild black hole has no ISCO, hence $r_{\rm isco}\to+\infty$.
-Equation (54) then confirms that $\delta$ indeed crosses zero.
-Consequently, under the isothermal constraint, the function $r_{\rm isco}(r_{h})$ is non-monotonic.
-(ii) RN-AdS ($q\neq 0$)
-Since the RN-AdS black hole exhibits a van der Waals–like phase transition, one has
+$\begin{cases}\delta\bigl(3r_{h,\min},\,r_{h,\min},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h,\min},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (56)
+where $r_{h,\min}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h,\min})=0$.
+For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h,\min}$.
+For $d>4$ and $P=0$, the metric (23) reduces to the higher-dimensional Schwarzschild solution, which admits no ISCO.
+This corresponds to $\lim_{P\to 0}r_{\rm isco}=+\infty$, see the end of Appendix E for a brief explanation.
+Equation (56) shows that $\delta>0$ at small $r_{h}$, and together with $\delta<0$ at sufficiently large $r_{h}$ this guarantees that $\delta$ crosses zero, implying that $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic on its full domain.
+(ii) RN–AdS ($q\neq 0$)
+Since the RN–AdS black hole exhibits a van der Waals–like phase transition, one has
 $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\leq 0\quad\text{for}\,\,T\geq T_{c},$ (57)
-where the equality holds precisely at the critical point $(T_{c},P_{c},r_{hc})$.
-According to the criterion (38), when $r_{h}\geq r_{h0}$ (where $r_{h0}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically.
-For $r_{h}<r_{h0}$, however, $\delta$ may become positive.
+where the equality holds precisely at the critical point $(r_{h,c},T_{c})$.
+According to the criterion (38), when $r_{h}\geq r_{h,e_{1}}$ (where $r_{h,e_{1}}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically.
+For $r_{h}<r_{h,e_{1}}$, however, $\delta$ may become positive.
 To determine this behavior, we examine the limit $r_{h}\to 0^{+}$.
@@
 $\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\!\left(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\right).$ (59)
-Meanwhile, from Appendix E, as $r_{h}\to 0^{+}$ one has
+Meanwhile, from Appendix F, as $r_{h}\to 0^{+}$ one has
 $V\,\partial_{M}\Phi_{1}\ \longrightarrow\ -\infty,\qquad\partial_{P}\Phi_{1}\ \longrightarrow\ 0.$ (60)
@@
 $\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\Bigl(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\Bigr)>0,\quad\Rightarrow\quad\frac{\mathrm{d}r_{\mathrm{isco}}}{\mathrm{d}r_{h}}<0.$ (61)
-This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic over the full domain of $r_{h}$.
+This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is nonmonotonic over the full domain of $r_{h}$.
 The same conclusion holds for $T<T_{c}$.
-Along both the small–black–hole and large–black–hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above.
+Along both the small–black hole and large–black hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above.
 The only difference is that for $T<T_{c}$ an intermediate spinodal (unstable) branch appears, on which $\mathrm{d}P/\mathrm{d}r_{h}>0$.
-Across this segment, the sign of $\delta$ may reverse, thereby altering the local monotonicity of $r_{\mathrm{isco}}(r_{h})$.
-In all cases, however, under the isothermal constraint $r_{\mathrm{isco}}(r_{h})$ is globally non-monotonic.
-Figure 6 presents the numerically computed profiles of $r_{\mathrm{isco}}(r_{h})$.
+Along this segment $\delta$ could, in principle, change sign, although this appears unlikely and the branch is dynamically skipped by the first-order phase transition.
+In any case, viewed globally, $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic for every fixed $T_{0}$.
+These results again confirm our previous analysis that $r_{\rm isco}(r_{h})$ can indeed exhibit nonmonotonic behavior in the isothermal ensemble.
+To verify the above analysis, we perform the numerical check, whose outcome is displayed in Fig.
+5.
 It can be seen that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for both charged and uncharged cases.
-The distinction lies in that, for the uncharged (Sch-AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ where $P(r_{h})=0$, whereas for the charged (RN-AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
-Appendix F likewise shows that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{isco}}(r_{h})$.
-Consequently, $P(r_{\mathrm{isco}})$ preserves the full local SS–shaped structure of $P(r_{h})$ across the phase transition.
-An explicit example for fixed parameters is shown in Fig.
-7.
+The distinction lies in that, for the uncharged (Sch–AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h,\min}=(d-3)/(4\pi T_{0})$, whereas for the charged (RN–AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
 Figure 5: Behavior of $r_{\rm isco}(r_{h})$ along isothermal paths in $d=4$.
+The breakdown of monotonicity naturally raises two questions.
+First, as noted above, when $r_{c}(r_{h})$ becomes nonmonotonic, can the jump $\Delta_{c}=r_{c}(r_{h,l})-r_{c}(r_{h,s})$ across the phase transition be reduced or even vanish?
+For RN–AdS black holes, the answer is no.
+In Appendix G, we prove that the extremal radius $r_{h,e}$ of $r_{\rm ps}(r_{h})$ always lies to the left of the small- and large-black hole radii, $r_{h,s}$ and $r_{h,l}$, namely $r_{h,e}<r_{h,s}<r_{h,l}$.
+So the phase transition always occurs on the increasing branch, indicating that $r_{\rm ps}$ undergoes a jump together with $r_{h}$ at the phase transition.
+This behavior is clearly illustrated in Figs.
+6 and 7, where the $P(r_{c})$ curve preserves the characteristic S-shaped structure of $P(r_{h})$.
+But whether this conclusion holds for other black holes must be examined case by case.
+If instead $r_{h,s}<r_{h,e}<r_{h,l}$, the values of $r_{\rm ps}(r_{h,s})$ and $r_{\rm ps}(r_{h,l})$ could become very close or even identical, potentially weakening or even destroying the reliability of using the circular orbit radius to encode the phase transition.
+It is therefore essential to analyze, from a theoretical perspective, to what extent phase transitions can be encoded in black hole images.
+This is precisely the role of the general criterion (21), which allows one to diagnose the monotonicity of $r_{\rm ps}(r_{h})$ in an arbitrary black hole spacetime and thereby assess the reliability of using circular orbit radii to encode phase transitions.
+Moreover, it is worth noting that, although the proof in Appendix G is presented for the RN–AdS black hole, the underlying method is general and can be applied to other black holes.
+In particular, we find that the extremal radius of $r_{\rm isco}(r_{h})$ always lies below that of $r_{\rm ps}(r_{h})$.
+We expect this ordering to persist in other spacetimes.
+Second, can nonmonotonic behavior encode additional thermodynamic information in black hole images?
+Given that both the black hole shadow and accretion-disk images are closely tied to $r_{\rm ps}$, we argue that the answer is positive.
+A simple idea is that the nonmonotonic evolution of black hole images may distinguish isobaric from isothermal ensembles.
+Further possibilities will be explored in our future work.
+Figure 6: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$.
+The solid and dashed lines correspond to $P(r_{h})$ and $P(r_{\rm ps})$, respectively.
+Both sets of lines are plotted for three temperatures: $T=0.8T_{c}$ (red), $T=T_{c}$ (green), and $T=1.2T_{c}$ (blue).
 Figure 7: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm isco})$ for $d=4$ and $q=1$.
-It is worth noting that, according to the first law of thermodynamics, a non-monotonic $M(r_{h})$ implies that the same mass $M$ can correspond to two distinct pressures $P$.
-Since both $r_{\mathrm{isco}}$ and $r_{h}$ depend on $P$, one might expect $r_{\mathrm{isco}}(r_{h})$ to be less prone to non-monotonic behavior.
-However, our results show that non-monotonicity indeed occurs, even for the Sch-AdS black hole.
-To better understand this, we examine the criterion (38).
-Because $\partial_{M}\Phi_{1}<0$ and $\partial_{P}\Phi_{1}>0$, $r_{\mathrm{isco}}(r_{h})$ is guaranteed to be monotonic as long as ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ have opposite signs.
-Yet the first law,
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (60)
-requires that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ share the same sign in certain ranges of $r_{h}$.
-For the $d$-dimensional RN-AdS black hole, ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is non-monotonic and can take negative values.
-Since the first law demands ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, the two derivatives become of the same sign, allowing the criterion (38) to cross zero.
-For the $d$-dimensional Sch-AdS black hole, comparing
-$T_{0}S^{\prime}(r_{h})=\frac{4\pi}{c_{d}}\,(d-2)\,T_{0}\,r_{h}^{\,d-3},$ (61)
-with
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{2(d-3)\,r_{h}^{\,d-4}}{(d-1)c_{d}}+\frac{4\pi(d-2)T_{0}\,r_{h}^{\,d-3}}{(d-1)c_{d}},$ (62)
-immediately reveals their relative size in the small-horizon regime. For instance, at the minimal value $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, one finds
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{d}{d-1}\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3},$ (63)
-$T_{0}S^{\prime}(r_{h})$ $=(d-2)\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3}.$ (64)
-It is clear that
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<T_{0}S^{\prime}(r_{h}),$
-which, through the first law $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ requires $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}>0$.
-Thus ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ acquire the same sign, allowing the criterion (38) to cross zero.
+The solid and dashed lines correspond to $P(r_{h})$ and $P(r_{\rm isco})$, respectively.
+Both sets of lines are plotted for three temperatures: $T=0.8T_{c}$ (red), $T=T_{c}$ (green), and $T=1.2T_{c}$ (blue).
 IV Critical-exponent invariance
-Critical exponents quantify the universal scaling near the black-hole critical point.
-They are conventionally defined with the horizon radius $r_{h}$, yet the interchangeability of $r_{\rm c}$ with $r_{h}$ suggests that one may equally define them using the circular-orbit radius $r_{\rm c}$.
-In this section we prove that, for both isobaric and isothermal constraints, the mapping $r_{\rm c}(r_{h})$ is locally monotonic at criticality, so the replacement $r_{h}\mapsto r_{\rm c}$ amounts to a smooth linear reparametrization in a neighborhood of the critical point.
-Consequently, critical exponent defined with $r_{\rm c}$ coincides with its counterpart defined with $r_{h}$.
-The order-parameter exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
+In the previous section, we showed that the mapping $r_{c}(r_{h})$ can exhibit global nonmonotonicity.
+On the one hand, this implies that encoding phase transitions in the circular orbit radius is not always reliable.
+On the other hand, from a thermodynamic viewpoint it raises a deeper issue: Whether such global nonmonotonicity can affect the universal critical behavior governed by the neighborhood of the critical point.
+Specifically, near criticality, the distinction between the coexisting phases becomes progressively weaker and vanishes exactly at the critical point.
+In conventional thermodynamics, such critical behavior is found to be universal and can be quantitatively characterized by critical exponents.
+For example, the classical Van der Waals system exhibits the order-parameter critical exponent $\beta=1/2$, which has also been shown to hold for many AdS black holes when the thermodynamic volume or the horizon radius is adopted to define the order parameter.
+Recently, Wei et al.
+demonstrated that in the isobaric ensemble the four-dimensional RN–AdS black hole yields the same value $\beta=1/2$ when the photon sphere radius is used to define the order parameter, and their numerical results for higher dimensional RN–AdS black holes remain very close to $1/2$ [19].
+These findings motivate us to test whether such universal critical behavior persists in different ensembles, especially when the mapping $r_{c}(r_{h})$ becomes globally nonmonotonic.
+In this section we prove that, under both isobaric and isothermal constraints, the mapping $r_{c}(r_{h})$ is necessarily locally monotonic at criticality.
+As a result, the reparametrization $r_{h}\mapsto r_{c}$ reduces to a smooth linear transformation in a neighborhood of the critical point.
+This guarantees that the critical exponents defined using $r_{c}$ coincide exactly with those obtained from $r_{h}$, so the universal critical behavior is fully preserved.
+The order-parameter critical exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
 $\Delta r_{h}:=r_{h,l}-r_{h,s}\sim|t|^{\beta},\qquad t:=\frac{T-T_{c}}{T_{c}}\to 0^{-},$ (62)
-with the mean-field value $\beta=1/2$. We will show that the circular-orbit radius inherits the same scaling, namely
-$\Delta r_{c}\;:=\;r_{c}\!\big(r_{h,l}\big)-r_{c}\!\big(r_{h,s}\big)\;\sim\;|t|^{1/2}.$ (66)
-We first perform a Taylor expansion of $P(r_{h},T)$ about the critical point $(r_{h,c},T_{c})$ up to the first nonvanishing orders allowed by the criticality conditions $\bigl.(\partial_{r_{h}}P)\bigr|_{c}=0$ and $\bigl.(\partial_{r_{h}}^{2}P)\bigr|_{c}=0$:
-$P(r_{h},T)-P_{c}$ $=\left.(\partial_{T}P)\right|_{c}\,\Delta T+\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T\,\Delta r$ (67)
-$+\frac{1}{2}\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)^{2}$
-$+\frac{1}{6}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{3}$
-$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
-where $\Delta r:=r_{h}-r_{h,c}$ and $\Delta T:=T-T_{c}$. At fixed temperature $T$, differentiating with respect to $r_{h}$ gives
-$\partial_{r_{h}}P(r_{h},T)$ $=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T$ (68)
-$+\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)$
-$+\frac{1}{2}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{2}$
-$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
-On the coexistence line for $T<T_{c}$, the Eq. (65) gives
-$\Delta r=(r_{h,l}-r_{h,c})-(r_{h,s}-r_{h,c})=\Delta r_{h}\sim|t|^{1/2},$ (69)
-so each power counts as $\Delta r^{n}\sim|t|^{n/2}$. Substituting this into (67) and keeping the leading orders in $|t|$ yields
-$\partial_{r_{h}}P(r_{h},T)$ (70)
-$=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,T_{c}\,t+\frac{1}{2}\left.(\partial_{r_{h}}^{3}P)\right|_{c}\,|t|+\mathcal{O}\!\big(|t|^{3/2}\big),$
-Thus $\bigl(\partial P/\partial r_{h}\bigr)_{T}=\mathcal{O}(|t|)\to 0$ as $t\to 0^{-}$. Along an isotherm, our criterion
-$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right)$ (71)
-combines with the first law
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ (72)
-to give, as $t\to 0^{-}$,
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{c}\,S^{\prime}(r_{h,c})+\mathcal{O}(|t|),\qquad\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=\mathcal{O}(|t|).$ (73)
-Since $S^{\prime}(r_{h,c})>0$, $V>0$, and $\Gamma_{\epsilon}\neq 0$, $\partial_{M}\Phi_{\epsilon}$ and $\partial_{P}\Phi_{\epsilon}$ are finite at the critical point, we obtain the fixed-sign leading term
-$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{T_{c}\,S^{\prime}(r_{h,c})}{\Gamma_{\epsilon}}\,\partial_{M}\Phi_{\epsilon}\ +\ \mathcal{O}(|t|),$ (74)
-so $r_{c}(r_{h})$ is strictly monotonic for sufficiently small $|t|$. (In the photon case, $\partial_{P}\Phi_{0}=0$ further collapses the subleading $\mathcal{O}(|t|)$ contribution.)
-Monotonicity (74) implies the first-order linearization at $r_{h,c}$,
-$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\bigl(|r_{h}-r_{h,c}|\bigr),$ (75)
-where $C:=\left.{dr_{c}}/{dr_{h}}\right|_{c}\neq 0$. Therefore on the coexistence line
-$\Delta r_{c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (76)
-and the critical exponent is unchanged:
-$\beta_{(r_{c})}=\beta_{(r_{h})}=\tfrac{1}{2}\quad\text{(mean-field)}.$ (77)
-For the isobaric process, $r_{c}(r_{h})$ is globally monotonic, so the above conclusion applies verbatim.
-In summary, under either isobaric or isothermal constraints, the reparameterization $r_{h}\!\mapsto\!r_{c}$ does not change the critical exponents.
-This is consistent with the numerical findings of Wei et al.
-in the isobaric photon-sphere case [12], for which we have provided an analytic proof.
-V Discussion
-In this work, we derived the general criterion (21), applicable to any parameter path in spherically symmetric black holes, which clarifies how circular orbits encode thermodynamic behavior.
-The criterion reveals that the monotonic relation between the circular-orbit radius $r_{\mathrm{c}}$ and the horizon radius $r_{h}$ is influenced by two factors: the intrinsic properties of the black hole and its circular orbits, and the parameter path chosen in the extended thermodynamic space.
-While the former never breaks monotonicity, the latter plays a decisive role.
-Motivated by this, we investigated two thermodynamically significant paths–the isobaric and isothermal processes–using the $d$-dimensional RN-AdS black hole as a test case.
-Our analysis demonstrates that the first law of thermodynamics directly constrains the monotonicity of $r_{\mathrm{c}}(r_{h})$.
-Under the isobaric constraint, the first law enforces $\mathrm{d}M/\mathrm{d}r_{h}\neq 0$, ensuring that $M(r_{h})$ and hence $r_{\mathrm{c}}(r_{h})$ remain monotonic.
-In contrast, under the isothermal constraint the first law allows $\mathrm{d}M/\mathrm{d}r_{h}=0$, which decouples the extrema of $M(r_{h})$ and $P(r_{h})$ so that a single mass $M$ corresponds to two distinct pressures $P$.
-This inevitably makes $r_{\mathrm{ps}}(r_{h})$ non-monotonic, since the photon-sphere radius $r_{\mathrm{ps}}$ is $P$-independent whereas $r_{h}$ depends on $P$.
-An exception occurs in the Sch-AdS limit ($q=0$), where $M(r_{h})$ remains monotonic; the emergence of electric charge $q$ thus breaks this monotonicity and, in turn, that of $r_{\mathrm{ps}}(r_{h})$.
-Interestingly, for the ISCO case, both $r_{\mathrm{isco}}$ and $r_{h}$ depend explicitly on $P$, which might seem to favor monotonicity.
-However, for the RN-AdS family we find that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for all $d$, including the Sch-AdS limit.
-This arises because the first law requires $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ to share the same sign in certain domains, making the prefactor in the criterion (38) change sign and thus breaking monotonicity.
-Overall, our results indicate that $r_{\mathrm{c}}(r_{h})$ tends to be monotonic under isobaric evolution but non-monotonic under isothermal evolution.
-It is worth noting that this non-monotonicity does not spoil the monotonic behavior of $r_{\rm c}(r_{h})$ in the local phase-transition region, nor does it affect its behavior near the critical point.
-As a result, the jump in $r_{\rm c}$ remains a reliable signature of the transition, and the critical exponents defined through $r_{\rm c}$ coincide with those obtained from $r_{h}$.
-These features are closely tied to the black hole first law and are expected to extend to more general classes of black holes.
-These results fill an important gap in the study of the non-monotonic behavior of $r_{\mathrm{c}}(r_{h})$ and provide a more robust foundation for identifying black hole phase transitions through optical observables.
-They also offer an additional avenue for probing black hole thermodynamics: the monotonicity (or lack thereof) of $r_{\mathrm{c}}(r_{h})$ can be used to distinguish whether a black hole evolves along an isobaric or an isothermal process.
-This idea is physically well motivated.
-By the entropy-increase law, the horizon radius $r_{h}$ must grow with time, and if $r_{\mathrm{c}}(r_{h})$ is non-monotonic, the resulting optical appearance of the black hole will exhibit correspondingly distinct behavior.
-Moreover, the influence of black hole thermodynamics on the monotonicity of $r_{\mathrm{c}}(r_{h})$ is itself a noteworthy feature deserving further investigation.
+We will show that the circular orbit radius inherits the same scaling, namely
+$\Delta r_{c}\;:=\;r_{c}\big(r_{h,l}\big)-r_{c}\big(r_{h,s}\big)\;\sim\;|t|^{\beta}.$ (63)
+The proof of Eq.
+(63) is straightforward.
+We first show that $r_{\rm c}(r_{h})$ is monotonic in the neighborhood of the critical point and then demonstrate that such a monotonic reparametrization leaves the critical exponents unchanged.
+Near the critical point one has ${\rm d}P/{\rm d}r_{h}\!\to\!0$, so that the isothermal criterion (38) reduces to
+$\frac{{\rm d}r_{\rm c}}{{\rm d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\frac{{\rm d}M}{{\rm d}r_{h}}\,\partial_{M}\Phi_{\epsilon},$ (64)
+which coincides exactly with the isobaric criterion (27).
+Consequently, in both ensembles the monotonic behavior of $r_{\rm c}(r_{h})$ near the critical point is entirely determined by that of $M(r_{h})$.
+In the isobaric ensemble, the first law enforces ${\rm d}M/{\rm d}r_{h}>0$, and $r_{\rm c}(r_{h})$ remains monotonic.
+In the isothermal ensemble, the first law gives
+$\frac{{\rm d}M}{{\rm d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{{\rm d}P}{{\rm d}r_{h}},$ (65)
+which simplifies near the critical point to
+$\frac{{\rm d}M}{{\rm d}r_{h}}=T_{0}\,S^{\prime}(r_{h}),$ (66)
+again ensuring ${\rm d}M/{\rm d}r_{h}>0$ and hence the monotonicity of $r_{\rm c}(r_{h})$. Therefore, in both the isobaric and isothermal ensembles,
+$C:=\left.\frac{{\rm d}r_{\rm c}}{{\rm d}r_{h}}\right|_{c}\neq 0.$ (67)
+We next expand $r_{\rm c}(r_{h})$ about the critical radius $r_{h,c}$ as
+$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\left(|r_{h}-r_{h,c}|\right).$ (68)
+This immediately yields
+$\Delta r_{\rm c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (69)
+which shows that, provided $C\neq 0$, the leading behavior of $\Delta r_{\rm c}$ near the critical point is governed by $\Delta r_{h}$. As a result, the critical exponent remains unchanged:
+$\beta_{(r_{c})}=\beta_{(r_{h})}.$ (70)
+This result is noteworthy: Although $r_{\rm c}(r_{h})$ may exhibit global nonmonotonic behavior, it must be monotonic in the neighborhood of the critical point, which ensures that the critical exponent remains invariant under reparametrization.
+This shows that the circular orbit radius encodes not only the existence of phase transitions but also their universal critical properties.
+Moreover, it is worth emphasizing that this proof does not rely on the specific form of the metric (23), but only on the criterion (38) and the first law.
+The result therefore applies, in principle, to any black hole with $\partial_{M}\Phi_{\epsilon}\neq 0$, which points to a more fundamental connection between black hole gravitational geometry and thermodynamics.
+V Summary and Discussion
+In this paper, we derived the monotonicity criterion (21) that applies to arbitrary thermodynamic paths in the parameter space of any static, spherically symmetric black hole, thereby clarifying when the mapping between the horizon radius $r_{h}$ and the circular orbit radius $r_{\rm c}$ is monotonic.
+In this framework, the monotonicity of $r_{\rm c}(r_{h})$ is governed by two factors: (i) the intrinsic geometric/orbital properties of the black hole and (ii) the chosen path in the extended thermodynamic parameter space.
+For any nonextremal black hole admitting an unstable photon sphere or an ISCO, the first factor does not by itself induce nonmonotonicity, whereas the second is decisive.
+Motivated by this, we take the dd–dimensional RN–AdS black hole as a concrete example and analyze two thermodynamically cases: the isobaric and isothermal ensembles.
+Our results show that the monotonicity of $r_{\rm c}(r_{h})$ is completely different in the two ensembles, and this contrast is closely tied to the first law of black hole thermodynamics.
+In the isobaric ensemble, the sign of $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ is determined by $\partial_{M}\Phi_{\epsilon}$ together with $\mathrm{d}M/\mathrm{d}r_{h}$.
+For RN–AdS, $\partial_{M}\Phi_{\epsilon}\neq 0$, while the first law enforces ${\rm d}M/{\rm d}r_{h}>0$, so ${\rm d}r_{\rm c}/{\rm d}r_{h}$ does not change sign and $r_{\rm c}(r_{h})$ is globally monotonic (both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ are strictly increasing).
+This monotonic increase guarantees that, whenever a phase transition occurs, $r_{\rm c}$ jumps together with $r_{h}$, thereby ensuring that the circular orbit radius reliably encodes the phase transition information.
+By the universality of the first law, this conclusion is not restricted to RN–AdS but extends to any black hole with $\partial_{M}\Phi_{\epsilon}\neq 0$.
+In contrast, in the isothermal ensemble the sign of $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ is controlled by four factors: $\partial_{M}\Phi_{\epsilon}$, $\partial_{P}\Phi_{\epsilon}$, $\mathrm{d}M/\mathrm{d}r_{h}$, and $\mathrm{d}P/\mathrm{d}r_{h}$.
+For the photon sphere one has $\partial_{P}\Phi_{0}=0$, so the sign of $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ again depends only on $\partial_{M}\Phi_{0}$ and $\mathrm{d}M/\mathrm{d}r_{h}$, but now the first law allows $\mathrm{d}M/\mathrm{d}r_{h}$ to cross zero.
+As a result, $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ may change sign and $r_{\rm ps}(r_{h})$ can become nonmonotonic.
+This is precisely what happens for the RN–AdS black hole, whose $r_{\rm ps}(r_{h})$ indeed exhibits nonmonotonic behavior, whereas the Sch–AdS case is exceptional: The absence of charge eliminates any zero of $\mathrm{d}M/\mathrm{d}r_{h}$, so $r_{\rm ps}(r_{h})$ remains monotonic.
+For the ISCO, one has $\partial_{P}\Phi_{1}>0$, so the sign of $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ is jointly controlled by $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$.
+The first law forces these two derivatives to share the same sign in certain ranges of $r_{h}$, thereby making sign changes in $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ possible.
+This possibility is realized explicitly in RN–AdS and Sch–AdS, where $r_{\rm isco}(r_{h})$ is fully nonmonotonic.
+Given the universality of the first law, we expect similar nonmonotonic behavior to appear for other black holes evolving along isotherms.
+In principle, such nonmonotonicity could reduce the jump in $r_{\rm c}$ at the phase transition, or even eliminate it, thereby undermining the reliability of using $r_{\rm c}$ to encode the phase transition.
+Fortunately, we have shown that, for RN–AdS, the phase transition always takes place on an increasing branch of $r_{\rm c}(r_{h})$, so the reliability of using $r_{\rm c}$ to encode the phase transition is not compromised.
+Whether this remains true for other black holes requires further case-by-case checks: One must first establish whether $r_{\rm c}(r_{h})$ becomes nonmonotonic in the isothermal ensemble and then determine whether this nonmonotonicity alters the jump amplitude $\Delta r_{\rm c}=r_{{\rm c},l}-r_{{\rm c},s}$ across the phase transition.
+Both steps rest on the monotonicity criterion (21), so in this sense our work provides an effective theoretical framework for assessing the observability of black hole phase transitions.
+It is also worth emphasizing that the nonmonotonic behavior of $r_{\rm c}(r_{h})$ in the isothermal ensemble has, to our knowledge, gone unnoticed so far.
+Thus, previous studies have largely focused on extracting phase transition information from black hole optical appearances, whereas we now identify an additional channel: The nonmonotonic evolution of the optical appearance can provide a diagnostic of the underlying thermodynamic ensemble.
+Since any nonmonotonicity of $r_{\rm c}(r_{h})$ should be reflected in the black hole image, such a test is feasible in principle.
+Of course, from an observational perspective, the relevant observables are the shadow and the accretion–disk image rather than $r_{\rm ps}$ or $r_{\rm isco}$ directly.
+Thus our next step is to analyze how the shadow and accretion–disk images evolve under isobaric versus isothermal processes.
+Finally, we prove that whenever a black hole exhibits critical behavior, the first law alone guarantees invariance of the critical exponents under the reparametrization $r_{h}\!\to r_{\rm c}$ for any spacetime satisfying $\partial_{M}\Phi_{\epsilon}\neq 0$.
+This shows that circular particle orbits not only encode phase transitions but also retain the universal critical behavior of black hole thermodynamics.
+This generality hints at a deeper geometric–thermodynamic correspondence that merits further study.
 Acknowledgements.
-This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
+We would like to thank Zhongzhinan Dong and Jiawei Chen for helpful discussions. This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
 Appendix A Preservation of Functional Behavior under Monotonic Reparameterization
@@
 To demonstrate that the nature of the stationary points is preserved, we differentiate once more:
-$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=F^{\prime\prime}\bigl(x(y)\bigr)\,[x^{\prime}(y)]^{2}+F^{\prime}\bigl(x(y)\bigr)\,x^{\prime\prime}(y).$ (82)
-At a stationary point $y_{m}$ where $F^{\prime}\bigl(x(y_{m})\bigr)=0$, this reduces to
-$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=F^{\prime\prime}(x_{m})\,[x^{\prime}(y_{m})]^{2}.$ (83)
-Since $x^{\prime}(y_{m})]^{2}>0$, it follows that
-$F^{\prime\prime}(x_{m})>0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})>0,$ (84)
-$F^{\prime\prime}(x_{m})<0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})<0.$
+$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^{2}+\frac{\mathrm{d}F}{\mathrm{d}x}\frac{\mathrm{d}^{2}x}{\mathrm{d}y^{2}}.$ (75)
+At a stationary point $y_{m}$ where ${\mathrm{d}F}/{\mathrm{d}x}=0$, this reduces to
+$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}\left.\left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^{2}\right|_{y_{m}}.$ (76)
+Since $\left({\mathrm{d}x}/{\mathrm{d}y}\right)^{2}>0$, it follows that
+$\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}>0$ $\ \Longrightarrow\ \left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}>0,$ (77)
+$\left.\frac{\mathrm{d}^{2}F}{\mathrm{d}x^{2}}\right|_{x_{m}}<0$ $\ \Longrightarrow\ \left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}<0.$
 Hence, under any smooth and monotonic reparameterization $x\mapsto y$, a local minimum (maximum) of $F$ remains a local minimum (maximum) of $\widetilde{F}$.
@@
 If $F$ is identified with the thermodynamic temperature $T$ or pressure $P$ of a black hole, and the variables are smoothly and monotonically related as $x=r_{h}$ and $y=r_{c}$, then under the mapping $r_{h}\mapsto r_{c}$ the functions $T(r_{c})$ and $P(r_{c})$ retain their characteristic van der Waals–like $S$-shaped structure.
-As a result, during the small-to-large black-hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
-Appendix B Constructive counterexample in Sch-AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
-For the four-dimensional Sch-AdS black hole, the metric function reads
+As a result, during the small-to-large black hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
+Appendix B Constructive counterexample in Sch–AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
+For the four-dimensional Sch–AdS black hole, the metric function reads
 $f(r)=1-\frac{2M}{r}+\frac{8\pi P}{3}\,r^{2},$ (78)
-where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon-sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
+where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
 $r_{\rm ps}=3M,$ (79)
@@
 Figure 8 illustrates $r_{\rm ps}(r_{h})$ for nine representative values of $P$, without imposing any specific relation between $P$ and $M$.
-Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch-AdS black hole.
+Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch–AdS black hole.
 Colored solid curves denote nine different pressures $P$, and the black dashed curve shows the path-constrained case with $k<0$ and $b>0$.
@@
 which agrees with Eq. (83).
-Appendix C Local Existence of $M(r_{h})$ and $P(r_{h})$
+Appendix C Conditions of the photon sphere and ISCO
+For photons ($\epsilon=0$) the radial function (5) reduces to
+$\mathcal{R}(r)=E^{2}-\frac{f(r)\,L^{2}}{r^{2}}.$ (88)
+Imposing the circular-orbit condition, $\mathcal{R}^{\prime}(r_{c})=0$, gives
+$\mathcal{R}^{\prime}(r_{\rm ps})=-\frac{f^{\prime}(r_{\rm ps})\,L^{2}}{r_{\rm ps}^{2}}+\frac{2f(r_{\rm ps})\,L^{2}}{r_{\rm ps}^{3}}=0.$ (89)
+Eliminating $L^{2}$ yields
+$\Phi_{0}(r_{\rm ps}):=r_{\rm ps}\,f^{\prime}(r_{\rm ps})-2f(r_{\rm ps})=0,$ (90)
+which corresponds to the photon-sphere equation (7). Imposing the instability of the photon circular orbit, $\mathcal{R}^{\prime\prime}(r_{\rm ps})>0$, gives
+$\mathcal{R}^{\prime\prime}(r_{\rm ps})=-\frac{L^{2}}{r_{\rm ps}^{3}}\,\underbrace{\big[r_{\rm ps}f^{\prime\prime}(r_{\rm ps})-f^{\prime}(r_{\rm ps})\big]}_{\;\equiv\ \partial_{r}\Phi_{0}(r_{\rm ps})}>0.$ (91)
+Since $L^{2}/r_{\rm ps}^{3}>0$, it follows that
+$\Gamma_{0}:=\partial_{r}\Phi_{0}(r_{\rm ps})=r_{\rm ps}f^{\prime\prime}(r_{\rm ps})-f^{\prime}(r_{\rm ps})<0,$ (92)
+which corresponds to the stability criterion (9).
+For massive particles ($\epsilon=1$) the radial function (5) reduces to
+$\mathcal{R}(r)=E^{2}-f(r)\!\left(1+\frac{L^{2}}{r^{2}}\right).$ (93)
+The conditions
+$\mathcal{R}^{\prime}(r_{\rm isco})=\mathcal{R}^{\prime\prime}(r_{\rm isco})=0,$ (94)
+give two algebraic equations for $L$ and $r_{\rm isco}$:
+$\mathcal{R}^{\prime}(r_{\rm isco})=\frac{2f(r_{\rm isco})\,L^{2}}{r_{\rm isco}^{3}}-f^{\prime}(r_{\rm isco})\!\left(1+\frac{L^{2}}{r_{\rm isco}^{2}}\right)=0,$ (95)
+$\mathcal{R}^{\prime\prime}(r_{\rm isco})=-\frac{6f(r_{\rm isco})L^{2}}{r_{\rm isco}^{4}}$ $+\frac{4f^{\prime}(r_{\rm isco})L^{2}}{r_{\rm isco}^{3}}$ (96)
+$-f^{\prime\prime}(r_{\rm isco})\!\left(1+\frac{L^{2}}{r_{\rm isco}^{2}}\right)=0.$
+Eliminating $L^{2}$ yields
+$\mathcal{R}^{\prime\prime}(r_{\rm isco})=\frac{2\,\Phi_{1}(r_{\rm isco})}{r_{\rm isco}^{2}\,f^{\prime}(r_{\rm isco})-2\,r_{\rm isco}f(r_{\rm isco})}=0,$ (97)
+where we define
+$\Phi_{1}(r_{\rm isco})=r_{\rm isco}\,f(r_{\rm isco})\,f^{\prime\prime}(r_{\rm isco})$ $-2\,r_{\rm isco}\,[f^{\prime}(r_{\rm isco})]^{2}$ (98)
+$+3\,f(r_{\rm isco})\,f^{\prime}(r_{\rm isco})=0$
+which corresponds to the ISCO equation (11). To determine the sign of
+$\Gamma_{1}:=\partial_{r}\Phi_{1}(r_{\rm isco}),$ (99)
+we rewrite Eq. (97) as
+$\mathcal{R}^{\prime\prime}(r_{\rm c})=\frac{2\,\Phi_{1}(r_{\rm c})}{\,r_{\rm c}^{2}f^{\prime}(r_{\rm c})-2r_{\rm c}f(r_{\rm c})\,}$, (100)
+where $r_{\rm c}$ denotes the radius of a timelike circular orbit in the neighborhood of the ISCO.
+By definition of the ISCO, orbits with $r_{\rm c}>r_{\rm isco}$ are stable $\bigl(\mathcal{R}^{\prime\prime}(r_{\rm c})<0\bigr)$, whereas those with $r_{\rm c}<r_{\rm isco}$ are unstable $\bigl(\mathcal{R}^{\prime\prime}(r_{\rm c})>0\bigr)$.
+Next, We analyze the sign of $\mathcal{R}^{\prime\prime}(r_{\rm c})$, which is fixed by the numerator and denominator of the ratio above.
+We abbreviate the denominator as
+$D(r_{\rm c}):=r_{\rm c}^{2}f^{\prime}(r_{\rm c})-2r_{\rm c}f(r_{\rm c})=r_{\rm c}^{4}\,\partial_{r}\!\left(\frac{f}{r^{2}}\right)\Big|_{r_{\rm c}},$ (101)
+where $V(r)\propto f(r)/r^{2}$ is the photon effective potential.
+For $r>r_{\rm ps}$ one has $\partial_{r}V<0$, and circular orbits near the ISCO satisfy $r_{\rm c}>r_{\rm ps}$.
+Therefore $D(r_{\rm c})<0$.
+We now examine the numerator.
+Expanding $\Phi_{1}$ about the ISCO,
+$\Phi_{1}(r_{\rm c})=\Phi_{1}(r_{\rm isco})+\partial_{r}\Phi_{1}(r_{\rm isco})\,(r_{\rm c}-r_{\rm isco})+O\bigl[(r_{\rm c}-r_{\rm isco})^{2}\bigr],$ (102)
+and using $\Phi_{1}(r_{\rm isco})=0$, we obtain near $r_{\rm isco}$:
+$\mathrm{sgn}\left[\mathcal{R}^{\prime\prime}(r_{\rm c})\right]=\mathrm{sgn}\!\left[\frac{\partial_{r}\Phi_{1}(r_{\rm isco})\,(r_{\rm c}-r_{\rm isco})}{D(r_{\rm isco})}\right].$ (103)
+Consequently, to ensure $\mathcal{R}^{\prime\prime}(r_{\rm c})>0$ for $r_{\rm c}<r_{\rm isco}$ and $\mathcal{R}^{\prime\prime}(r_{\rm c})<0$ for $r_{\rm c}>r_{\rm isco}$, the slope of $\Phi_{1}$ at the ISCO must be positive, $\Gamma_{1}:=\partial_{r}\Phi_{1}(r_{\rm isco})>0$, which is precisely the stability criterion (12).
+Appendix D Local Existence of $M(r_{h})$ and $P(r_{h})$
 Consider the two constraint equations at the horizon,
@@
 $\Delta=-\,\frac{4c_{d}}{d-2}\,r_{h}^{\,4-d}\neq 0.$ (108)
-Appendix D Proof of the Positivity of $\partial_{P}\Phi_{1}$
+Appendix E Proof of the Positivity of $\partial_{P}\Phi_{1}$
 $\partial_{P}\Phi_{1}=\frac{16\pi\,r_{\mathrm{isco}}}{(d-2)(d-1)}$ [ $\Bigl[8+4(d-1)(d-2)q^{2}r_{\mathrm{isco}}^{6-2d}$ (109)
@@
 Hence,
-$\mathrm{sgn}\,\partial_{P}\Phi_{1}=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (105)
+$\mathrm{sgn}\left(\partial_{P}\Phi_{1}\right)=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},+\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (114)
 To determine the sign of $\partial_{P}\Phi_{1}$, we first examine the sign of $\Delta_{x}$. Under the isothermal constraint one has
@@
 $\Phi_{1}(r,M,q,P,d)=P\,\partial_{P}\Phi_{1}+\widetilde{\Phi}_{1}(r,M,q,d).$ (121)
-Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
+Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to+\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
 $\partial_{P}\Phi_{1}\to 0.$ (122)
-This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
+This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to+\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
 $\Phi_{1}=\frac{2Mr_{\mathrm{isco}}^{2}(r_{\mathrm{isco}}-6M)+18Mq^{2}r_{\mathrm{isco}}-8q^{4}}{r_{\mathrm{isco}}^{5}}.$ (123)
@@
 From Eq.
-(104), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
+(113), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{\rm min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
 Therefore, for both the Schwarzschild and RN black holes one always has $x\!\mid_{P=0}>x_{+}$, and consequently $\partial_{P}\Phi_{1}>0$ in this limit.
@@
 Since $\Gamma_{1}>0$, this requires $r_{\mathrm{isco}}$ to decrease, meaning that $x$ moves toward $x_{+}$.
-As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to\infty}$.
+As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to+\infty}$.
 Consequently, for any finite $P$, the corresponding $x$ must lie within the interval $(x_{+},\,x\!\mid_{P=0})$, which guarantees $\partial_{P}\Phi_{1}>0$ throughout.
@@
 For higher dimensions ($d>4$), the argument becomes even simpler.
-At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to\infty$).
+At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to+\infty$).
 As $P$ increases, $r_{\mathrm{isco}}$ moves inward from infinity toward $x_{+}$, again ensuring $\partial_{P}\Phi_{1}>0$.
@@
 $\partial_{P}\Phi_{1}>0.$ (124)
-Appendix E Asymptotic Relations in the Small-Horizon Limit
+The limit used above,
+$\lim_{P\to 0}r_{\rm isco}=+\infty,$ (125)
+can be argued as follows. For the higher-dimensional Schwarzschild black hole ($q=0$, $d>4$), the ISCO condition takes the form
+$\Phi_{1}(r_{\rm isco})$ $=\frac{16\pi P\,r_{\rm isco}}{(d-2)(d-1)}\Bigl[8-c_{d}(d^{2}-1)M\,r_{\rm isco}^{3-d}\Bigr]$ (126)
+$\quad-c_{d}(d-3)M\,r_{\rm isco}^{2(1-d)}\Bigl[\,c_{d}(d-1)M\,r_{\rm isco}^{3}$
+$\qquad\qquad\qquad\qquad\qquad+(d-5)\,r_{\rm isco}^{d}\Bigr].$
+If $r_{\rm isco}$ remains finite as $P\to 0$, the first (pressure-dependent) term vanishes and $\Phi_{1}$ reduces to
+$\Phi_{1}(r_{\rm isco})=-c_{d}(d-3)M\,r_{\rm isco}^{2(1-d)}$ $\Bigl[\,c_{d}(d-1)M\,r_{\rm isco}^{3}$ (127)
+$+(d-5)\,r_{\rm isco}^{d}\Bigr],$
+which admits no positive root for $r_{\rm isco}$ when $d>4$.
+Thus a finite $r_{\rm isco}$ is incompatible with the ISCO condition in this limit.
+By contrast, if $r_{\rm isco}\to+\infty$ as $P\to 0$, the second term in $\Phi_{1}$ is suppressed, and one finds
+$\Phi_{1}(r_{\rm isco})\;\simeq\;\frac{128\pi P\,r_{\rm isco}}{(d-2)(d-1)},$ (128)
+so that the product $P\,r_{\rm isco}$ can tend to zero as $P\to 0$. Therefore the ISCO condition can be satisfied only if
+$\lim_{P\to 0}r_{\rm isco}=+\infty.$ (129)
+For the higher-dimensional RN black hole ($q\neq 0$, $d>4$), the ISCO condition takes the form
+$\Phi_{1}(r_{\rm isco})=$ $-2(d-3)\,q\,r_{\rm isco}^{\,2(1-2d)}\bigl(-q\,r_{\rm isco}^{3}+r_{\rm isco}^{d}\bigr)^{2}$ (130)
+$\times\bigl[2(d-2)\,q\,r_{\rm isco}^{3}+(d-5)\,r_{\rm isco}^{d}\bigr]$
+$+\frac{32\pi P\,r_{\rm isco}}{(d-2)(d-1)}\Bigl\{4-(d-1)\,q\,r_{\rm isco}^{\,3-2d}$
+$\qquad\times\bigl[-2(d-2)\,q\,r_{\rm isco}^{3}+(d+1)\,r_{\rm isco}^{d}\bigr]\Bigr\},$
+where we have imposed the extremality condition $M=2q/c_{d}$. Repeating the same analysis as in the Schwarzschild case then yields Eq. (125).
+Appendix F Asymptotic Relations in the Small-Horizon Limit
 We now examine the asymptotic behavior of $V\,\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ in the small-horizon limit ($r_{h}\to 0^{+}$), focusing on their leading divergent terms.
-For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix D for details). The thermodynamic volume takes the simple form
+For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to+\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix E for details). The thermodynamic volume takes the simple form
 $V=\frac{\Omega_{d-2}\,r_{h}^{\,d-1}}{d-1},$ (131)
@@
 $M\sim\frac{1}{r_{h}^{\,d-3}},\qquad P\sim\frac{1}{r_{h}^{\,2(d-2)}}.$ (135)
-For the ISCO, the divergence $P\!\to\!\infty$ implies (see Appendix D)
+For the ISCO, the divergence $P\!\to\!+\infty$ implies (see Appendix E)
 $r_{\mathrm{isco}}^{\,d-3}\to x_{+}=\frac{(d-1)(d+1)c_{d}M+\sqrt{\Delta_{x}}}{16},$ (136)
@@
 $V\,\partial_{M}\Phi_{1}\sim-\frac{1}{r_{h}}\;\longrightarrow\;-\infty.$ (139)
-Appendix F Monotonicity of $r_{\rm c}(r_{h})$ in the Phase-Transition Region
-F.1 Photon sphere ($\epsilon=0$)
-Our proof consists of two steps.
-First, we demonstrate that $r_{h0}$ lies to the left of $r_{h,\min}$.
-Second, we show that $P(r_{h0})>P(r_{h,\max})$, where $r_{h,\min}$ and $r_{h,\max}$ denote the locations of the minimum and maximum of $P(r_{h})$, respectively.
-This procedure circumvents the need to determine the small-black-hole radius $r_{h,s}$ from Maxwell’s equal-area law while still guaranteeing that $r_{h0}$ lies to the left of $r_{h,s}$.
-Step 1
-From Eq.
-(41) it follows that ${\mathrm{d}P}/{\mathrm{d}r_{h}}=0$ implies ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$.
-Since $M(r_{h})$ first decreases and then increases, $r_{h0}$ must therefore lie to the left of both extrema, namely on the small-black-hole branch.
-Step 2
-To establish the result for arbitrary $q$, we eliminate $q$ by expressing $P(r_{h})$ and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ in reduced variables:
-$P(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (125)
-$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (126)
-where $\tilde{r}_{h}=r_{h}/r_{h,c}$ and $\tilde{T}_{0}=T_{0}/T_{c}$, with $r_{h,c}=\sqrt{6}\,q$ and $T_{c}=1/(3\sqrt{6}\,\pi q)$ denoting the critical values.
-Using these relations, we numerically determine $r_{h0}$ and $r_{h,\max}$, and plot $P(r_{h0})$ (red curve) and $P(r_{h,\max})$ (blue curve) as functions of $\tilde{T}$ in Fig.
+Appendix G Monotonicity of $r_{\rm c}(r_{h})$ in the Phase Transition Region
+We take the four-dimensional case as an example and show that the phase transition occurs only along the increasing branch of $r_{\rm c}(r_{h})$.
+The proof strategy is simple: we numerically scan the parameter space, extract the extremal radius $r_{h,e}$ of $r_{\rm ps}(r_{h})$ and the small–black hole radius $r_{h,s}$ at phase coexistence, and then compare their values.
+In the isothermal ensemble the free parameters are the temperature $T_{0}$ and the charge $q$.
+Conveniently, the explicit qq–dependence can be removed by introducing reduced (dimensionless) variables
+$\tilde{r}_{h}=\frac{r_{h}}{r_{h,c}},\quad\tilde{M}=\frac{M}{M_{c}},\quad\tilde{P}=\frac{P}{P_{c}},\quad\tilde{T}=\frac{T}{T_{c}},$ (140)
+where the critical values at the thermodynamic critical point are
+$r_{h,c}$ $=\sqrt{6}\,Q,\qquad M_{c}=\frac{2\sqrt{6}\,Q}{3},$ (141)
+$P_{c}$ $=\frac{1}{96\pi\,Q^{2}},\qquad T_{c}=\frac{1}{3\sqrt{6}\,\pi\,Q}.$
+They are obtained by imposing the criticality conditions
+$\frac{\partial P}{\partial r_{h}}=\frac{\partial^{2}P}{\partial r_{h}^{2}}=0,\qquad\frac{\partial T}{\partial r_{h}}=\frac{\partial^{2}T}{\partial r_{h}^{2}}=0.$ (142)
+Using these reduced variables, we further obtain
+$\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (143)
+$\frac{\mathrm{d}\tilde{M}}{\mathrm{d}\tilde{r}_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (144)
+which no longer contain $q$ explicitly, so the conclusions derived from them apply to arbitrary charge.
+The only remaining control parameter is $\tilde{T}_{0}$.
+To ensure $\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})\geq 0$ for all $\tilde{r}_{h}$, one must have $\tilde{T}_{0}\geq\sqrt{2}/2$, while the first–order phase transition occurs only for $T_{0}<T_{c}$, i.e., $\tilde{T}_{0}<1$.
+Thus the relevant range is $\tilde{T}_{0}\in[\sqrt{2}/2,\,1)$.
+Within this interval, we numerically extract $r_{h,s}$ from $\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})$ via the Maxwell equal-area construction, and obtain $r_{h,e}$ from the zero of $\mathrm{d}\tilde{M}/\mathrm{d}\tilde{r}_{h}$.
+The resulting curves are shown in Fig.
 9.
-The black dashed vertical line at $\tilde{T}_{0}=\sqrt{2}/2$ marks the critical value above which $P>0$; the region $\tilde{T}_{0}<\sqrt{2}/2$ is therefore unphysical and excluded from consideration.
-From the remaining portion of the plot it is evident that $P(r_{h0})>P(r_{h,\max})$.
-Combined with the result of Step 1, this demonstrates that $r_{h0}$ lies to the left of $r_{h,s}$.
-To visualize this behavior, Fig.
-10 shows $P(r_{h})$ (red) and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ (blue) for several $\tilde{T}_{0}$.
-At $\tilde{T}_{0}=0.4$, $P(r_{h0})<P(r_{h,\max})$, but $P(r_{h})<0$ and is thus unphysical.
-When $\tilde{T}_{0}\geq\sqrt{2}/2$, $P(r_{h})\geq 0$ and $P(r_{h0})>P(r_{h,\max})$, with $r_{h,s}$ and $r_{h,l}$ clearly located to the right of $r_{h0}$.
-F.2 ISCO ($\epsilon=1$)
-With the photon-sphere case established, the ISCO analysis becomes straightforward. From the criterion
+From this figure, one clearly sees that $r_{h,e}<r_{h,s}$ for all $\tilde{T}_{0}$ in the allowed range.
+Equivalently, the zero of $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ always lies to the left of $r_{h,s}$, demonstrating that the phase transition takes place on the increasing branch of $r_{\rm ps}(r_{h})$.
+We now extend this conclusion to the ISCO case. From the criterion (38), i.e.,
 $\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right),$ (145)
-we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is momentarily ignored, the zeros of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ coincide with those of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{0}$, since both are controlled entirely by $\mathrm{d}M/\mathrm{d}r_{h}$.
-Because $\partial_{M}\Phi_{1}<0$ and $\mathrm{d}M/\mathrm{d}r_{h}$ increases monotonically, ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ is strictly decreasing.
-Restoring the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}<0$ can only shift the zero of $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}$further to the left.
-Since the zero of $\mathrm{d}M/\mathrm{d}r_{h}$ already lies to the left of $r_{h,s}$ (as shown earlier), its additional leftward shift necessarily keeps it on the small-black-hole side.
-Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN-AdS black hole is globally non-monotonic, it remains locally monotonic within the phase-transition region, thereby preserving the discontinuous jump across the transition.
-Figure 9: Plots of $p(r_{h0})$ (red) and $p(r_{h,\max})$ (blue) as functions of $\tilde{T}_{0}$. At $\tilde{T}_{0}=\sqrt{2}/2$, one has $p(r_{h})>0$ for all $r_{h}$.
-Figure 10: Behaviors of $P(r_{h})$ and $\mathrm{d}M/\mathrm{d}r_{h}$ for three representative values of $\tilde{T}_{0}$.
+we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is temporarily neglected, the zeros of $(\mathrm{d}M/\mathrm{d}r_{h})\,\partial_{M}\Phi_{1}$ coincide with those of $(\mathrm{d}M/\mathrm{d}r_{h})\,\partial_{M}\Phi_{0}$, since both are determined solely by $\mathrm{d}M/\mathrm{d}r_{h}$.
+Moreover, $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ is an increasing function of $r_{h}$ in a neighborhood of the zero.
+Restoring $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ then adds a positive contribution $(1/\Gamma_{1})\,\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}>0$, which lifts the curve and shifts the zero of $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ to smaller $r_{h}$.
+Consequently, the extremum of $r_{\rm isco}(r_{h})$ lies to the left of that of $r_{\rm ps}(r_{h})$, and hence also to the left of $r_{h,s}$, so the phase transition again occurs on the increasing branch of $r_{\rm isco}(r_{h})$.
+Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN–AdS black hole is globally non-monotonic, it remains locally monotonic within the phase transition region, thereby preserving the discontinuous jump across the phase transition.
+Figure 9: Plots of $\tilde{r}_{h,e}$, $\tilde{r}_{h,s}$, and $\tilde{r}_{h,l}$ as functions of $\tilde{T}_{0}$.
+The curves $\tilde{r}_{h,s}$ and $\tilde{r}_{h,l}$ intersect at point A, where $r_{h,s}=r_{h,l}=r_{h,c}$.